### v1

# On the gcd graphs over polynomial rings and related topics

Abstract.
Gcd-graphs over the ring of integers modulo $n$ are a natural generalization of unitary Cayley graphs. The study of these graphs has foundations in various mathematical fields, including number theory, ring theory, and representation theory. Using the theory of Ramanujan sums, it is known that these gcd-graphs have integral spectra; i.e., all their eigenvalues are integers. In this work, inspired by the analogy between number fields and function fields, we define and study gcd-graphs over polynomial rings with coefficients in finite fields. We establish some fundamental properties of these graphs, emphasizing their analogy to their counterparts over $\operatorname{\mathbb{Z}}.$
Key words and phrases:
Cayley graphs, Function fields, Ramanujan sums, Symmetric algebras, Integral graphs, Prime graphs, Perfect graphs.
2020 Mathematics Subject Classification:
Primary 05C25, 05C50, 05C51
JM is partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant R0370A01. He gratefully acknowledges the Western University Faculty of Science Distinguished Professorship 2020-2021. TTN is partially supported by an AMS-Simons Travel Grant. Parts of this work were done while he was a postdoc associate at Western University. He thanks Western University for their hospitality. NDT is partially supported by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.04-2023.21
1. Introduction
Let $n$ be a positive integer. The unitary Cayley graph on the ring of integers modulo $n$ is defined as the Cayley graph $G_{n}={\rm Cay}(\operatorname{\mathbb{Z}}/n,U_{n})$ where $\operatorname{\mathbb{Z}}/n$ is the ring of integers modulo $n$ and $U_{n}=(\operatorname{\mathbb{Z}}/n)^{\times}$ its unit group. More specifically, $G_{n}$ is equipped with the following data.
  1. (1)
    The vertex set of $G_{n}$ is $\operatorname{\mathbb{Z}}/n$.
  2. (2)
    Two vertices $a,b\in V(G_{n})$ are adjacent if and only if $a-b\in U_{n}$.
The unitary Cayley graph $G_{n}$ was first formally introduced in [18] even though we can trace it back to the work of Evans and Erdős [11]. Due to its elegance and simplicity, the unitary Cayley graph has been further studied and generalized by various works in the literature. For example, [1] studies the unitary Cayley graph of a finite commutative ring. [15] generalizes this study further to finite rings which are not necessarily commutative. In [26], a subset of the authors studies various arithmetic and graph-theoretic properties of the $p$-unitary Cayley graph as defined by [27, 28] (when $p=1$, these graphs are precisely unitary Cayley graphs). We refer the interested readers to [3, 8, 18, 26] and the references therein for some further topics in this line of research.
As explained in [18, Section 4], a particularly intriguing arithmetical property of $G_{n}$ is that its spectrum can be described by the theory of circulant matrices and Ramanujan sums. A consequence of this fact is that the unitary Cayley graph has an integral spectrum; i.e., all of its eigenvalues are integers. In [18], the authors note that the unitary Cayley graph is not the sole graph exhibiting an integral spectrum. They identify a closely related family of graphs sharing this characteristic, known as gcd-graphs which we now recall. Let $D=\{d_{1},d_{2},\ldots,d_{k}\}$ be a set of proper divisors of $n$. The gcd-graph $G_{n}(D)$ is the graph with the following data.
  1. (1)
    The vertex set of $G_{n}(D)$ is $\operatorname{\mathbb{Z}}/n.$
  2. (2)
    Two vertices $a,b\in G_{n}$ are adjacent if and only if $\gcd(a-b,n)\in D.$
We remark that the unitary Cayley graph $G_{n}$ is nothing but $G_{n}(\{1\}).$ It is known that the spectrum of $G_{n}(D)$ is a summation of various Ramanujan sums (see [18, Section 4]). In particular, all of its eigenvalues are integers. It turns out that, the converse is also true.
Theorem 1.1.
(See [29, Theorem 7.1]) A $\operatorname{\mathbb{Z}}/n$-circulant graph $G$ is an integral graph if and only if $G=G_{n}(D)$ for some $D.$
Let $\operatorname{\mathbb{F}}_{q}$ be a finite field. As observed by Andre Weil in [30], there is a strong analogy between the ring $\operatorname{\mathbb{Z}}$ of integers and the ring $\operatorname{\mathbb{F}}_{q}[x]$ of polynomials with coefficients in $\operatorname{\mathbb{F}}_{q}$ (and more generally, between number fields and function fields). Consequently, it is of reasonable interest to define and investigate gcd-graphs over $\mathbb{F}_{q}[x]$. Fortunately, the analogous definition of gcd-graphs over $\mathbb{F}_{q}[x]$ is relatively straightforward. More specifically, let $f\in\operatorname{\mathbb{F}}_{q}[x]$ be a non-zero element in $\operatorname{\mathbb{F}}_{q}[x]$. Let $D=\{f_{1},f_{2},\ldots,f_{k}\}$ be a subset of the set of divisors $\operatorname{{\rm Div}}(f)$ of $f$; i.e, $f_{i}\mid f$ for all $1\leq i\leq k.$ Let
$S_{D}=\{g\in\operatorname{\mathbb{F}}_{q}[x]/f|\gcd(g,f)\in D\}.$
Definition 1.2.
We denote by $G_{f}(D)$ the gcd-Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S_{D}).$ More precisely, it is the graph equipped with the following data
  1. (1)
    The vertex set of $G_{f}(D)$ is the finite ring $\operatorname{\mathbb{F}}_{q}[x]/f.$
  2. (2)
    Two vertices $u,v$ are adjacent if and only if $u-v\in S_{D}$. In other words, $\gcd(u-v,f)\in D.$
Remark 1.3.
In general, the greatest common divisor is only defined up to associates. In our case, however, we can make a canonical choice for the great common divisor by requiring it to be a monic polynomial. Therefore, unless we explicitly state, we will assume throughout this article that all involved polynomials are monic.
Example 1.4.
Fig. 1 shows the graph $G_{f}(D)$ where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$ and $D=\{x,x+1\}$. It is a regular graph of degree $4$
Figure 1. The gcd-Cayley graph $G_{x(x+1)}(\{x,x+1\})$
1.1. Outline
In this article, we study some foundational properties of gcd-graphs defined over $\mathbb{F}_{q}[x]$. In Section 2, we recall some standard definitions in graph theory that will be used throughout the article. Section 3 investigates the question of when a gcd-graph is connected and anti-connected. In Section 4, we provide the necessary and sufficient conditions for a gcd-graph to be bipartite. Section 5 explores another graph-theoretic property of gcd-graphs: their prime-property. Although we are not able to provide a complete answer, we present some rather general conditions under which this property holds. Section 6 examines the spectra of gcd-graphs using symmetric algebras and Ramanujan sums. We find it quite remarkable that the explicit formulas for these spectra are identical to those that appeared in the number field case. In Section 7, we provide some sufficient conditions for gcd-graphs to be perfect. Finally, in Section 8, we investigate whether a given graph can be realized as an induced subgraph of a gcd-graph. As a by-product, we prove an analogous result to a theorem of Erdős and Evans in the number field case.
1.2. Code
The code that we develop to generate gcd-graphs and do experiments on them can be found at [23].
2. Background from graph theory
In this section, we recall some basic concepts in graph theory that we will use throughout this article.
Definition 2.1 (The complete graph $K_{n}$).
$K_{n}$ is the graph on $n$ vertices which are pairwisely adjacent.
Definition 2.2 (Tensor product of graphs).
Let $G,H$ be two graphs. The tensor product $G\times H$ of $G$ and $H$ (also known as the direct product) is the graph with the following data:
  1. (1)
    The vertex set of $G\times H$ is the Cartesian product $V(G)\times V(H)$,
  2. (2)
    Two vertices $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G\times H$ if and only if $(g,g^{\prime})\in E(G)$ and $(h,h^{\prime})\in E(H).$
Definition 2.3 (Wreath product).
Let $G,H$ be two graphs. We define the wreath product of $G$ and $H$ as the graph $G*H$ with the following data
  1. (1)
    The vertex set of $G*H$ is the Cartesian product $V(G)\times V(H)$,
  2. (2)
    $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G*H$ if either $(g,g^{\prime})\in E(G)$ or $g=g^{\prime}$ and $(h,h^{\prime})\in E(H)$.
Definition 2.4 (Graph morphism).
Let $G_{1}$ and $G_{2}$ be two graphs. We define a graph morphism between $G_{1}$ and $G_{2}$ to be a map from $V(G_{1})$ to $V(G_{2})$ $f$ that preserves edges. More precisely, if $u,v\in V(G_{1})$ are adjacent in $G_{1}$, then $f(u),f(v)$ are adjacent in $G_{2}$.
Definition 2.5 (Induced subgraph).
Let $G_{1}$ and $G_{2}$ be two graphs. We say that $G_{1}$ is an induced subgraph of $G_{2}$ if there exists a graph morphism $f:G_{1}\to G_{2}$ such that the following conditions hold.
  1. (1)
    The map $f\colon V(G_{1})\to V(G_{2})$ is injective.
  2. (2)
    For $u,v\in G_{1}$, $u,v$ are adjacent if and only if $f(u)$ and $f(v)$ are adjacent in $G_{2}.$
3. Connectedness of gcd graphs
In this section, we will give the necessary and sufficient conditions for $G_{f}(D)$ to be connected. Unlike the number field case where these conditions are rather straightforward, the function field case is a bit more complicated. This is partially due to the fact that the unitary Cayley graph on $\operatorname{\mathbb{F}}_{q}[x]/f$ is not always connected. We start our discussion with the following simple observation.
Lemma 3.1.
If $G_{f}(D)$ is connected then $\gcd(f_{1},\ldots,f_{k})=1$.
Proof.
Let $f_{0}=\gcd(f_{1},\ldots,f_{k})$. We observe that if two vertices $u$ and $v$ are adjacent then $f_{0}$ divides $u-v$. Now since $G_{f}(D)$ is connected, there is a path connecting 00 and $1$. This implies that $f_{0}$ divides $1-0=1$. Hence $\gcd(f_{1},\ldots,f_{k})=1$. ∎
In the case of $\operatorname{\mathbb{Z}}$, the converse of 3.1 is true as well; i.e. the gcd-graph $G_{n}(D)$, where $D=\{d_{1},d_{2},\ldots,d_{k}\}$ is connected if and only if ${\rm gcd}(d_{1},d_{2},\ldots,d_{k})=1.$ In the case of $\operatorname{\mathbb{F}}_{q}[x]$, this condition is not sufficient. For example, let $D=\{1\}$, $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2},f=x(x+1).$ Then $\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{2}\times\operatorname{\mathbb{F}}_{2}.$ In this case
$G_{f}(\{1\})\cong K_{2}\times K_{2},$
which is not connected (in general, this is the only obstruction where the unitary graph over a commutative ring fails to be connected, see [8] and 3.3)). In the case of the gcd-graphs, we have the following result.
Theorem 3.2.
$G_{f}(D)$ is connected if the following conditions hold.
  1. (1)
    $\gcd(f_{1},f_{2},\ldots,f_{k})=1$.
  2. (2)
    The unitary Cayley graph $G_{f}(\{1\})$ is connected.
Proof.
We need to show that $R=\langle S_{D}\rangle$ the abelian group generated by $S_{D}$. Let $a\in R=\operatorname{\mathbb{F}}_{q}[x]/f.$ Because $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, we can find $a_{1},a_{2},\ldots,a_{k}\in R$ such that
$a=\sum_{i=1}^{k}a_{i}f_{i}.$
Since $G_{f}(\{1\})$ is connected, for each $1\leq i\leq d$, we can find write
$a_{i}=\sum_{j=1}^{n_{i}}m_{ij}s_{ij},$
where $m_{ij}\in\operatorname{\mathbb{Z}}$ and $s_{ij}\in R^{\times}.$ Consequently, we can write
$a=\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}m_{ij}s_{ij}f_{i}.$
By definition, $\gcd(s_{ij}f_{i},f)=f_{i}\in D$. This shows that $a\in\langle S_{D}\rangle$. Since this is true for all $a$, we conclude that $R=\langle S_{D}\rangle$. ∎
We can classify all $f$ such that $G_{f}(\{1\})$ is not connected.
Proposition 3.3.
(See [8, Lemma 4.33]) $G_{f}(\{1\})$ is not connected if and only if $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$
By 3.2 and 3.3, we have the following corollary.
Corollary 3.4.
Suppose that one of the following conditions holds.
  1. (1)
    $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$.
  2. (2)
    $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\nmid f.$
Then $G_{f}(D)$ is connected if and only if $\gcd(f_{1},f_{2},\ldots,f_{k})=1.$
We now deal with the case where $G_{f}(\{1\})$ is not connected. By 3.3, this implies that $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$
Lemma 3.5.
Let $g$ be a divisor of $f$. Then for every polynomial $h\in\operatorname{\mathbb{F}}_{q}[x]$
$\gcd(h,g)=\gcd(\gcd(h,f),g).$
Proof.
Let $m=\gcd(h,f).$ We need to show that $\gcd(h,g)=\gcd(m,g).$ We first claim that $\gcd(h,g)\mid\gcd(m,g).$ By definition, $\gcd(h,g)\mid g.$ We also have $\gcd(h,g)\mid\gcd(h,f)=m.$ Therefore, $\gcd(h,g)\mid\gcd(m,g).$
Conversely, we claim that $\gcd(m,g)\mid\gcd(h,g).$ This is clear since $m\mid h$.
In summary, we have $\gcd(m,g)\mid\gcd(h,g)$ and $\gcd(h,g)\mid\gcd(m,g).$ This shows that $\gcd(m,g)=\gcd(h,g).$
Proposition 3.6.
Let $g$ be a divisor of $f$ and $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g$ be the canonical projection map. Let
$\overline{D}=\{\gcd(f_{i},g)|1\leq i\leq k\}.$
Then $\Phi_{f,g}(S_{D})\subseteq S_{\overline{D}}$. Consequently, $\Phi_{f,g}\colon G_{f}(D)\to G_{g}(\overline{D})$ is a graph morphism.
Proof.
Suppose that $a\in S_{D}.$ Then, there exists $i$ such that $\gcd(a,f)=f_{i}.$ By 3.5, we know that
$\gcd(a,g)=\gcd(\gcd(a,f),g)=\gcd(f_{i},g).$
This shows that $a\in S_{\overline{D}}.$ We conclude that $\Phi_{f,g}(S_{D})\subseteq S_{\overline{D}}$. ∎
Corollary 3.7.
If $G_{f}(D)$ is connected then $G_{g}(\overline{D})$ is connected as well.
Proof.
Let $\bar{a},\bar{b}$ be two vertices in $G_{g}(\overline{D}).$ Since $\Phi_{f,g}$ is surjective, we can find $a,b\in\operatorname{\mathbb{F}}_{q}[x]/f$ such that $\Phi_{f,g}(a)=\bar{a},\Phi_{f,g}(b)=\bar{b}.$ Since $G_{f}(D)$ is connected, there is a path $P$ from $a$ to $b.$ By 3.6, $\Phi_{f,g}(P)$ is a path from $\bar{a}$ to $\bar{b}.$ This shows that $G_{g}(\overline{D})$ is connected as well. ∎
We are now ready to state our theorem in the case $G_{f}(\{1\})$ is not connected.
Theorem 3.8.
Suppose that $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$ Let $\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/(x(x+1))$ be the canonical projection map and $\bar{D}$ as described above. Then $G_{f}(D)$ is connected if and only if
  1. (1)
    $\gcd(d_{1},d_{2},\ldots,d_{k})=1$, and
  2. (2)
    the graph $G_{x(x+1)}(\bar{D})$ is connected. This condition is equivalent to $|\overline{D}\setminus\{x(x+1)\}|\geq 2.$
Proof.
By 3.7, we know that if $G_{f}(D)$ is connected then $(1)$ and $(2)$ holds. Conversley, let us assume that $(1)$ and $(2)$ both hold. We will show that $G_{f}(D)$ is connected. This is equivalent to showing that $R:=\operatorname{\mathbb{F}}_{q}[x]/f=\langle S_{D}\rangle.$
The key idea of this proof is similar to the proof for 3.2. The main difficulty is to deal with the fact that $G_{f}(\{1\})$ is not connected in this case. We will do this step by step.
We first claim that if $g\in R$ such that $x(x+1)\mid g$ then $g=s_{1}+s_{2}$ where $s_{1},s_{2}\in R^{\times}.$ For an element $s\in R$, $s$ is a unit if and only $\Phi^{\text{ss}}(s)\in(R^{\text{ss}})^{\times}$ where $R^{\text{ss}}=R/J(R)$ is the semisimplification of $R$ (see [8, Proposition 4.30]). Therefore, for this statement, we can assume that $R=R^{\text{ss}}$; namely $f$ is a squarefree polynomial. If we write $f=x(x+1)f_{1}$ where $\gcd(x(x+1),f_{1})=1$ then we have the isomorphism
$R\cong\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))\times\operatorname{\mathbb{F}}_{2}[x]/f_{1}.$
Under this isomorphism, $g$ is sent to $(0,g_{1})$ where $g_{1}\in\operatorname{\mathbb{F}}_{2}[x]/f_{1}.$ Since $F_{2}[x]/f_{1}$ is a product of fields of order bigger than $2$, every element in it can be written as the sum of two units; say $g_{1}=t_{1}+t_{2}$ where $t_{1},t_{2}\in(\operatorname{\mathbb{F}}_{2}[x]/f)^{\times}.$ We then see that $g=s_{1}+s_{2}$ where $s_{1}=(1,t_{1})$ and $s_{2}=(1,t_{2})$. By definition $s_{1},s_{2}\in R^{\times}.$
We now claim that if $g\in R$ such that $x(x+1)\mid g$ then $g\in\langle S_{D}\rangle.$ Since $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, we can find $a_{1},a_{2},\ldots,a_{k}\in R$ such that
$a_{1}f_{1}+a_{2}f_{2}+\cdots+a_{k}f_{k}=1.$
Multiplying both sides with $g$, we get $g=\sum_{i=1}^{k}a_{i}gf_{i}.$ Since $x(x+1)\mid a_{i}g$, we can write $a_{i}g=s_{1i}+s_{2i}$ where $s_{1i},s_{2i}\in R^{\times}.$ This shows that
(3.1) $g=\sum_{i=1}^{k}(s_{1i}f_{i}+s_{2i}f_{i}).$
Since $s_{1},s_{2}\in R^{\times}$, $s_{1i}f_{i},s_{2i}f_{i}\in S_{D}.$ This shows that $g\in\langle S_{D}\rangle.$
Finally, let $g$ now be an arbitrary element in $R$. We claim that $g\in\langle S_{D}\rangle.$ By our assumption, the graph $G_{x(x+1)}(S_{\bar{D}})$ is connected, 00 and $\Phi_{f,x(x+1)}(g)$ are connected by a path. Consequently, we can write
$g\equiv\sum_{i}n_{i}\gcd(h_{i},x(x+1))\pmod{x(x+1)},$
where $n_{i}\in\operatorname{\mathbb{Z}}$ and $\gcd(h_{i},f)\in D$. We can check that over $\operatorname{\mathbb{F}}_{2}[x]$
$\gcd(h,x(x+1))\equiv h\pmod{x(x+1)},$
for all $h\in\operatorname{\mathbb{F}}_{2}[x].$ Therefore, we can write
$g\equiv\sum_{i}n_{i}h_{i}\pmod{x(x+1)}.$
By the previous case, we know that $g-\sum_{i=1}^{k}n_{i}h_{i}\in\langle S_{D}\rangle.$ This shows that $g\in\langle S_{D}\rangle$ as well. Since this is true for all $g$, we conclude that $R=\langle S_{D}\rangle.$
Remark 3.9.
The proof of 3.8 also implies that for the case $\operatorname{\mathbb{F}}_{2}[x]$ and $x(x+1)\mid f$, the map $\Phi_{f,g}:\operatorname{\mathbb{F}}_{2}[x]/f\to\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))$ has the property that $a$ and $b$ belong the same connected component in $G_{f}(D)$ if and only $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ belong to the same connected component in $G_{x(x+1)}(\overline{D}).$ In fact, by 3.7, if $a$ and $b$ are connected by a path then $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path. Conversely, if $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path in $G_{x(x+1)}(\overline{D})$, we can write
$a-b\equiv\sum_{i}n_{i}h_{i}\pmod{x(x+1)}.$
where $n_{i}\in\operatorname{\mathbb{Z}}$ and $\gcd(h_{1},f)\in D$. Because $x(x+1)R\subseteq\langle S_{D}\rangle$, we conclude that that $a-b\in\langle S_{D}\rangle.$ By definition, $a$ and $b$ are connected by a path in $G_{f}(D)$.
We conclude that $G_{f}(D)$ and $G_{x(x+1)}(\overline{D})$ has the same number of connected component. In particular, the number of connected component in $G_{f}(D)$ is at most $2$
Remark 3.10.
While most of our discussions in this section concern the connectedness of $G_{f}(D)$, similar statements hold for the anti-connectedness of $G_{f}(D)$ as well. In fact, the complement of $G_{f}(D)$ is precisely $G_{f}(\operatorname{{\rm Div}}(f)\setminus(D\cup\{f\})$ where $\operatorname{{\rm Div}}(f)$ is the set of all proper divisors of $f.$
4. Bipartite property
A graph $G=(V,E)$ is called a bipartite graph if $V$ can be partitioned into two disjoint sets $V_{1},V_{2}$ such that for every edge $(u,v)\in E$ either $u\in V_{1},v\in V_{2}$ or $u\in V_{2},v\in V_{1}$. Bipartite graphs model various real-life situations such as job assignments, resource allocation, stable marriage matching, gene-disease association, and much more. Additionally, from a theoretical point of view, bipartite graphs often provide a good test for theorems and algorithms on graphs. We refer the reader to [2] for some further discussions on this topic.
The goal of this section is to classify all $G_{f}(D)$ which is bipartite. We remark that while our focus is on gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x]$, everything we discuss in this section applies to gcd-graphs over $\operatorname{\mathbb{Z}}$ as well (see 4.5). To the best of our knowledge, even over $\operatorname{\mathbb{Z}}$, this topic has not been explored in the literature. Therefore, in addition to studying this problem for its own merits, we also hope to address a gap in the literature.
We start with the following lemma.
Lemma 4.1.
Suppose that $G_{f}(D)$ is connected and bipartite.
  1. (1)
    Let
    $I=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 0\pmod{2}\right\},$
    $I_{1}=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 1\pmod{2}\right\},$
    Then $I$ is an subgroup of index $2$ in $(R,+).$ Furthermore, $I$ and $I_{1}$ are independent set such that $V(G_{f}(D))=I\bigsqcup I_{1}.$
  2. (2)
    If $G_{f}(\{1\})$ is connected, then $I$ is an ideal in $R$ as well.
  3. (3)
    If $G_{f}(\{1\})$ is not connected, which is equivalent to $x(x+1)\mid f$ and $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ by 3.3, then $x(x+1)R\subseteq I.$
Proof.
  1. (1)
    Let us consider the first part of this lemma. Because $G_{f}(D)$ is bipartite, we can write
    $V(G_{f}(D))=A\bigsqcup B$
    where $A,B$ are two independent sets in $G_{f}(D).$ Without loss of generality, we can assume that $0\in A$. We claim that $A=I$. By the proof of [4, Proposition 2.6], we know that $A$ is a subgroup of $(R,+)$ of index $2$ By definition, if $h_{i}\in S_{D}$, then $(0,h_{i})\in E(G_{f}(D))$. Because $A$ and $B$ are disjoint independent set in $G_{f}(D)$, $h_{i}\in B$. Furthermore, since $A$ is a subgroup of index $2$ in $R$, $h_{i}+h_{j}\in A$ for all $h_{i},h_{j}\in S_{D}.$ Consequently, we must have $I\subseteq A.$
    By our assumption that $G_{f}(D)$ is connected we know that $R=\langle S_{D}\rangle$ and hence $I\sqcup I_{1}=\langle S_{D}\rangle=R.$ Clearly if $a\in R\setminus I=I_{1}$ then $2a\in I$. From this, we can see that $I$ is a subgroup whose index is at most $2$ in $R$. Additionally, because $I\subseteq A$ and $A$ has index $2$ in $R$, we must have that $I=A$ and $I_{1}=B.$
  2. (2)
    Suppose that $G_{f}(\{1\})$ is connected. We will show that $I$ is an ideal in $R$ as well. The idea is similar to the proof of 3.2 so we will be brief. Specifically, we note that for each $s\in R^{\times}$, $sI=I.$ Since $\langle R^{\times}\rangle=R$, this shows that $aI\subseteq I$ for all $a\in R.$ We conclude that $I$ is an ideal in $R$.
  3. (3)
    Finally, part $(3)$ follows from Equation Eq. 3.1.
Corollary 4.2.
Suppose that either $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$ or $\gcd(x(x+1),f)=1$. Then $G_{f}(D)$ is not a bipartite graph.
Proof.
If $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$ or $\gcd(x(x+1),f)=1$, we know that $G_{f}(\{1\})$ is connected by 3.3. Furthermore, $\operatorname{\mathbb{F}}_{q}[x]/f$ has no ideal of index $2$ By 4.1, we conclude that $G_{f}(D)$ is not bipartite. ∎
We now consider the case $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $\gcd(x(x+1),f)\neq 1.$ We first consider the following easier case.
Theorem 4.3.
Suppose that $f$ is a polynomial in $\operatorname{\mathbb{F}}_{2}[x]$ such that $\gcd(x(x+1),f)\not\in\{1,x(x+1)\}.$ Then $G_{f}(D)$ is bipartite if and only if $\gcd(x(x+1),f)\nmid f_{i}$ for all $1\leq i\leq k.$
Proof.
Without loss of generality, we can assume that $x\mid f$ but $x+1\nmid f.$ First, let us assume that $x\nmid f_{i}$ for all $1\leq i\leq k$. In this case, we can see that $A=xR$ and $B=1+xR$ are independent subsets in $G_{f}(D)$ such that $V(G_{f}(D))=A\bigsqcup B.$ This shows that $G_{f}(D)$ is bipartite.
Conversely, let us assume that $G_{f}(D)$ is bipartite. We claim that $x\nmid f_{i}$ for all $1\leq i\leq k$. Since $x(x+1)\nmid f$, by 3.2 and 3.3, we know that $G_{f}(\{1\})$ is connected. Therefore, by 4.1, there exists an ideal $I$ of index $2$ in $R$ such that $I$ is an independent set in $G_{f}(D).$ We remark that since the only ideal of index $2$ in $R$ is $xR$, $I=xR$. Because $0\in I$ and $I$ is independent we must have
$xR\cap\{f_{1},f_{2},\ldots,f_{k}\}=\emptyset.$
In other words, $x\nmid f_{i}$ for all $1\leq i\leq k.$
Finally, let us consider the trickest case where $x(x+1)\mid f.$
Theorem 4.4.
Suppose that $x(x+1)\mid f$ and that $G_{f}(D)$ is connected. Let
$\overline{D}=\{\gcd(f_{i},x(x+1))|f_{i}\in D.\}$
Then $G_{f}(D)$ is bipartite if and only if $|\overline{D}|=2.$
Proof.
Suppose that $|\overline{D}|=2$. We claim that $G_{f}(D)$ is bipartite. Because $G_{f}(D)$ is connected, we know by 3.8 that $|\overline{D}\setminus\{x(x+1)\}|\geq 2.$ This shows that $\overline{D}\neq\{1,x(x+1)\}.$ We can check that $\overline{D}$ must be one of the following sets $\{1,x\},\{1,x+1\},\{x,x+1\}.$
First, we consider the case that $\overline{D}=\{1,x+1\}$. In this case $V(G_{f}(D))=xR\bigsqcup(1+xR)$ is a decomposition of $G_{f}(D)$ into a disjoint union of two independent sets. Similarly, if $\overline{D}=\{1,x\}$ then $V(G_{f}(D))=(x+1)R\bigsqcup(1+(x+1)R)$ is a decomposition of $G_{f}(D)$ into a disjoint union of two independent sets. Now, suppose that $\overline{D}=\{x,x+1\}.$ Let
$A=\{g\in R|g(0)=g(1)\},$
and
$B=\{g\in R|g(0)\neq g(1)\}=x+A.$
We can check that if $a_{1},a_{2}\in A$ then $\gcd(a_{1}-a_{2},x(x+1))\in\{1,x(x+1)\}.$ By definition, $(a_{1},a_{2})\not\in E(G_{f}(D)).$ This shows that $A$ is an independent set in $G_{f}(D)$. Similarly, $B$ is an independent set in $G_{f}(D)$ as well. We conclude that $G_{f}(D)$ is not bipartite.
Conversely, suppose $G_{f}(D)$ is bipartite. We claim that $|\overline{D}|=2.$ By 4.1, the subgroup
$I=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}n_{i}\equiv 0\pmod{2}\right\},$
is an independent set in $G_{f}(D).$ Furthermore, by part $(3)$ of 3.7, we know that $x(x+1)R\subset I.$ Because 00 is not connected to any node in $I$, we must have $x(x+1)R\cap S_{D}=\emptyset$. Consequently $x(x+1)\not\in\overline{D}.$ Suppose to the contrary that $|\overline{D}|\geq 3$. We then have $\overline{D}=\{1,x,x+1\}.$ This implies that $G_{x(x+1)}(\overline{D})$ is the complete graph $K_{4}.$ In $K_{4}$, there is a path of length $2$ from 00 to any other vertices. From this property, we conclude that for each $g\in R$, we can write
$g\equiv\gcd(h_{1},x(x+1))+\gcd(h_{2},x(x+1))\pmod{x(x+1)},$
for some $h_{1},h_{2}\in S_{D}$. We then conclude that $g-h_{1}-h_{2}\in x(x+1)R\subset I$. Since $h_{1}+h_{2}\in I$, we conclude that $g\in I$ as well. This shows that $I=R$, which is a contradiction. ∎
Remark 4.5.
By an almost identical argument, we can show that the gcd graph $G_{n}(D)$ over $\operatorname{\mathbb{Z}}$ with $D=\{d_{1},d_{2},\ldots,d_{k}\}$ and $\gcd(d_{1},d_{2},\ldots,d_{k})=1$, is bipartite if and only if $2\mid n$ and $2\nmid d_{i}$ for all $1\leq i\leq k.$
5. Primeness of gcd graphs
For a given graph $G$, a homogeneous set in $G$ is a set $X$ of vertices of $G$ such that every vertex in $V(G)\setminus X$ is adjacent to either all or none of the vertices in $X$. A homogenous set $X$ is said to be non-trivial if $2\leq X<|V(G)|$. The graph $G$ is said to be prime if it does not contain any non-trivial homogeneous sets.
The concept of a homogeneous set appears in various branches of mathematics (see [6, 8, 12, 25] for some concrete examples). One of our main motivations comes from the fact that homogeneous sets allow us to decompose a network into a multilevel network of smaller graphs. From both theoretical and computational perspectives, such a decomposition is crucial for understanding the dynamics of multilevel networks (see [5, 13, 17, 25]). In [8], in collaboration with some other graph theorists, we completely classify prime unitary Cayley graphs on finite commutative rings. In this section, we study the problem for the gcd graph $G_{f}(D).$ We remark that for a graph $G$, a connected component of $G$ (or of its complement) is a homogeneous set in $G$. Therefore, while studying the primeness of $G$, it is safe to assume that $G$ is connected and anti-connected. For this reason, we will assume throughout this section that $G_{f}(D)$ is both connected and anti-connected (we refer the reader to Section 3 for precise conditions for these properties to hold).
An important property of homogeneous sets is that they are preserved under a graph isomorphism. For this reason, we start our discussion with the following observation.
Proposition 5.1.
Let $a\in(\operatorname{\mathbb{F}}_{q}[t]/f)^{\times}.$ Let $m_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/f$ be the multiplication by $a$ map. Then $m_{a}$ induces an automorphism on $G_{f}(D).$
Proof.
Since $a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}$, $m_{a}$ is an automorphism of $(\operatorname{\mathbb{F}}_{q}[x]/f,+)$. Furthermore, $m_{a}$ preserves $S_{D}$; i.e $aS_{D}=S_{D}.$ As a result, $m_{a}$ is an automorphism of $G_{f}(D).$
Proposition 5.2.
Assume that $G_{f}(\{1\})$ is connected. Then, the following conditions are equivalent.
  1. (1)
    $G_{f}(D)$ is not a prime graph.
  2. (2)
    There exists a non-trivial ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/f$ such that $I$ is a homogeneous set in $G_{f}(D).$
Proof.
Clearly $(2)$ implies $(1)$. Let us now prove $(1)$ implies $(2)$. The proof that we discuss here is quite similar to the one that we gave for [8, Theorem 4.1] and 3.2. Because $G_{f}(D)$ is not prime, we can find a maximal homogenous set $H$ containing $0$ By [8, Theorem 3.4], we know that $H$ is a subgroup of $\operatorname{\mathbb{F}}_{q}[x]/f$. We claim that it is an ideal as well. By 5.1, we conclude that $aH=H$ for all $a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}.$ Furthermore, since $G_{f}(\{1\})$ is connected, $aH\subseteq H$ for all $a\in\operatorname{\mathbb{F}}_{q}[x]/f$ as well. This shows that $H$ is an ideal in $\operatorname{\mathbb{F}}_{q}[x]/f.$
By 5.2, in order to study the prime property of $G_{f}(D)$, it is essentially equivalent to classify $g$ such that the ideal generated by $g$ is a homogenous set in $G_{f}(D).$ In order to do so, we first introduce the following lemmas.
Lemma 5.3.
Let $a,b,c\in\operatorname{\mathbb{F}}_{q}[x]$ such that $\gcd(a,b)=\gcd(b,c)=m$ and $c\mid f.$ Then, there exists $t\in\operatorname{\mathbb{F}}_{q}[x]$ such that
$\gcd(a-bt,f)=c.$
Proof.
By replacing $a,b,c,f$ by $\frac{a}{m},\frac{b}{m},\frac{c}{m},\frac{f}{m}$, we can assume that $m=1.$ Because $\gcd(b,c)=1$, we can find $t_{1}$ such that $c\mid a-bt_{1}.$ Let us write $a-bt_{1}=a_{1}c.$ We will look for $t=t_{1}+ct_{2}$ such that the condition $\gcd(a-bt,f)=c$ holds. By our choices of $t_{1},t_{2}$, this is equivalent to
$c=\gcd(a-b(t_{1}+ct_{2}),f)=\gcd(a_{1}c-bt_{2}c,f)=c\gcd(a_{1}-bt_{2},\frac{f}{c}).$
We remark that the relation $a-bt_{1}=a_{1}c$ and the fact that $\gcd(a,b)=1$ imply that $\gcd(a_{1},b)=1.$ By the Chinese remainder theorem, we can find $t_{2}$ such that $\gcd(a_{1}-bt_{2},\frac{f}{c})=1.$
Lemma 5.4.
Let $g$ be a divisor of $f$. Let $I_{g}$ be the ideal in $R=\operatorname{\mathbb{F}}_{q}[x]/f$ generated by $g.$ The induced graph on $I_{g}$ is isomorphic to $G_{f/g}(D_{g})$ where
$D_{g}=\left\{\frac{f_{i}}{g}|f_{i}\in D,g\mid f_{i}\right\}.$
Proof.
Every element in $I_{g}$ can be written in the form $gm$ for a unique $m\in\operatorname{\mathbb{F}}_{q}[x]/(f/g)$. Therefore we have a natural map $I_{g}\to\operatorname{\mathbb{F}}_{q}[x]/(f/g)$ sending $gm\mapsto m.$ Furthermore, for two elements $ga,gb\in I_{g}$, we have
$\gcd(ga-gb,f)=g\gcd(a-b,\frac{f}{g}).$
Therefore, we see that $\gcd(ga-gb,f)\in D$ if and only $\gcd(a-b,\frac{f}{g})\in D_{g}.$ From this, we conclude that the induced graph on $I_{g}$ is naturally isomorphic to $G_{f/g}(D_{g})$. ∎
Theorem 5.5.
Let $g\mid f$ be a divisor of $f$ and $I$ the ideal in $\operatorname{\mathbb{F}}_{q}[x]/f$ generated by $g.$ Let
$D_{1}=\left\{f_{i}\in D\mid g\nmid f_{i}\right\},\quad D_{2}=\left\{f_{i}\in D|g\mid f_{i}\right\}.$
As in 3.6 let
$\overline{D_{1}}=\left\{\gcd(f_{i},g)\mid f_{i}\in D_{1}\right\},\quad\widetilde{D_{2}}=\left\{\frac{f_{i}}{g}\mid f_{i}\in D_{2}\right\}.$
Then, the following statements are equivalent.
  1. (1)
    $I$ is a homogeneous set in $G_{f}(D)$.
  2. (2)
    $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$ where $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g$ is the canonical projection map.
Furthermore, if one of the above equivalent conditions holds, $G_{f}(D)$ is isomorphic to the wreath product $G_{g}(D_{1})*G_{f/g}(\widetilde{D_{2}}).$
Proof.
First, let us show that $(2)$ implies $(1)$. In other words, suppose that $g$ satisfies the condition that $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$. We claim that the ideal $I$ generated by $g$ is a homogeneous set in $G_{f}(D).$ In fact, let $a\not\in I$ be an arbitrary element in $\operatorname{\mathbb{F}}_{q}[x]/f$ and suppose that $a$ is adjacent to an element in $I$. By a translation, we can assume that $a$ is adjacent to 00 in $G_{f}(D).$ We claim that $a$ is adjacent to all elements in $I$ as well. Let $gt$ be an element in $I$. We need to show that
$\gcd(a-gt,f)\in D.$
Since $(a,0)\in E(G_{f}(D))$ we know that $\gcd(a,f)\in D.$ Because $a\not\in I$, we know further that $\gcd(a,f)\in D_{1}$. Additionally, by 3.5, we must have $\gcd(a,g)=\gcd(\gcd(a,f),g)\in\overline{D_{1}}.$ Again, by 3.5, we have
$\gcd(\gcd(a-gt,f),g)=\gcd(a-gt,g)=\gcd(a,g)\in\overline{D_{1}}.$
Because $D_{1}=\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)$, this shows that $\gcd(a-gt,f)\in D_{1}$ and hence $\gcd(a-gt,f)\in D$ as required.
Conversely, we claim that $(1)$ implies $(2)$. Suppose that $I$ is a homogeneous set in $G_{f}(D).$ We need to show that $D_{1}=\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)$. By 3.6, we always have $D_{1}\subset\Phi_{f,g}^{-1}(D_{1})\cap\operatorname{{\rm Div}}(f)$. Therefore, it is sufficient to show that $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)\subset D_{1}.$ Let $h\in\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)$. By definition, there exists $f_{i}\in D_{1}$ such that
$\gcd(h,g)=\gcd(f_{i},g).$
By 5.3, we can find $t\in\operatorname{\mathbb{F}}_{q}[x]$ such that
$\gcd(h-gt,f)=f_{i}.$
This shows that $(h,gt)\in E(G_{f}(D)).$ Since $I$ is homogenous and $h\not\in I$, we conclude that $(h,0)\in E(G_{f}(D))$ as well. By definition, $\gcd(h,f)\in D$. Since $h\mid f$, we conclude that $h\in D$ and hence $h\in D_{1}.$
In general, it seems unclear how to check the conditions mentioned in 5.5 explicitly. We discuss here a particular case when this can be done.
Proposition 5.6.
Let $f,g,I$ be as in 5.5. Assume further that $f_{i}\nmid f_{j}$ for all $i\neq j$ and $\gcd(f_{1},f_{2},\ldots,f_{k})=1$. Then $I$ is a homogeneous set in $G_{f}(D)$ if and only if the following conditions hold.
  1. (1)
    For each $1\leq i\leq k$, $f_{i}\mid g$ for all $i$.
  2. (2)
    Furthermore,
    ${\rm rad}\left(\frac{g}{f_{i}}\right)={\rm rad}\left(\frac{f}{f_{i}}\right).$
In particular, if $f$ is squarefree then $f=g.$
Proof.
Let us first assume that $I$ is homogenous. Let $g_{i}=\gcd(f_{i},g).$ We claim that if $f_{i}\in D_{1}$ then $g_{i}=f_{i}$ and hence $f_{i}\mid g.$ In fact, we have $\gcd(g_{i},g)=g_{i}=\gcd(f_{i},g).$ This shows that $g_{i}\in\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$ (by 5.5). Since $f_{i}\nmid f_{j}$ for all $i\neq j$, we must have $f_{i}=g_{i}$ and hence $f_{i}\mid g.$ By our assumption, $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, and hence we must have $D_{1}\neq\emptyset$. If $D_{2}\neq\emptyset$ then for each $f_{1}\in D_{1}$ and $f_{2}\in D_{2}$, we have $f_{1}\mid g\mid f_{2}$ which is a contradiction. Therefore, we must have $D_{2}=\emptyset.$ In summary, we just prove that $f_{i}\mid g$ for all $1\leq i\leq k.$ We now show that
${\rm rad}\left(\frac{g}{f_{i}}\right)={\rm rad}\left(\frac{f}{f_{i}}\right).$
Suppose this is not the case. We can find a non-constant irreducible polynomial $h$ of such that $\gcd(h,\frac{g}{f_{i}})=1$ and $h|\frac{f}{f_{i}}.$ We then see that $\gcd(hf_{i},g)=f_{i}=\gcd(f_{i},g).$ This implies that $hf_{i}\in D$. Since $f_{i}\mid hf_{i}$ and $f_{i}\neq hf_{i}$, this leads to a contradiction.
Let us now show the converse. By 5.5, we need to show that if $h\in\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)$ then $h\in D_{1}.$ Since $h\in\Phi_{f,g}^{-1}(\overline{D_{1}})$ we can find $f_{i}\in D_{1}$ such that $\gcd(h,g)=\gcd(f_{i},g)=f_{i}.$ Let us write $h=f_{i}h_{1}$ with $\gcd(h_{1},\frac{g}{f_{i}})=1.$ By our assumption that ${\rm rad}\left(\frac{g}{f_{i}}\right)={\rm rad}\left(\frac{f}{f_{i}}\right)$, we must have $\gcd(h_{1},\frac{f}{f_{i}})=1$ as well. Since $h_{1}\mid\frac{f}{f_{i}}$, we must have $h_{1}=1$ and hence $h=f_{i}.$
Corollary 5.7.
Let that $f$ be a squarefree polynomial. Suppose that $G_{f}(\{1\})$ is connected. Let $D=\{f_{1},f_{2},\ldots,f_{k}\}$ be a subset of $\operatorname{{\rm Div}}(f)$ such that the following conditions hold.
  1. (1)
    $f_{i}\nmid f_{j}$ for all $i\neq j.$
  2. (2)
    $G_{f}(D)$ is connected and anti-connected.
Then $G_{f}(D)$ is prime.
Proof.
Suppose that $G_{f}(D)$ is not prime. By 5.2, there exists a proper divisor $g\mid f$ such that the ideal $I$ generated by $g$ is homogeneous. By 5.6, we must have $g=f$ which is a contradiction. ∎
Remark 5.8.
The proof of 5.6 relies crucially on the divisibility relationship between $f_{i}$ and $f_{j}.$ It seems important to study this relationship systematically. We are studying this problem in a work in progress (see [22]).
Remark 5.9.
While we primarily focus on gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x]$ in this section, most statements have a straightforward analog for gcd-graphs over $\operatorname{\mathbb{Z}}.$
6. Spectrum of gcd graphs
The spectrum of gcd-graphs over $\operatorname{\mathbb{Z}}$ is described by the theory of Ramanujan sums (see [18, Section 4]), which in turn is a special case of Gauss sums (see [21]). As explained in [7], these sums are precisely values of Fekete polynomials at certain $n$-roots of unity. One might wonder whether such a similar statement for the spectrum of gcd-graphs holds in the context of function fields. While we are not able to find an analog of Fekete polynomials over function fields, the theory of Ramanujan sums does have an interesting analogy as we will explain in this section. This, however, is sufficient for us to describe explicitly the spectrum of gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x].$
6.1. Symmetric algebras
A key point in the theory of $\operatorname{\mathbb{Z}}/n$-circulant graph is the fact that the character group of $\operatorname{\mathbb{Z}}/n$ is isomorphic to $\operatorname{\mathbb{Z}}/n$:
$\operatorname{\mathbb{Z}}/n\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times}).$
This isomorphism can be obtained as follows. Fix a primitive $n$-root of unity $\zeta_{n}$ in $\operatorname{\mathbb{C}}$. Let $\chi_{1}\colon\operatorname{\mathbb{Z}}/n\to\operatorname{\mathbb{C}}^{\times}$ be the character defined by $\chi_{1}(m)=\zeta_{n}^{m}$ for all $m\in\operatorname{\mathbb{Z}}/n.$ For each $a\in\operatorname{\mathbb{Z}}/n$, let $\chi_{a}=\chi_{1}^{a}$ be the character of $\operatorname{\mathbb{Z}}/n$ defined by $\chi_{a}(b)=\zeta_{n}^{ab}=\chi_{1}^{a}(b).$ The following proposition is standard.
Proposition 6.1.
The map $a\mapsto\chi_{1}^{a}$ gives an isomorphism between $\operatorname{\mathbb{Z}}/n$ and $\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times}).$
In summary, once we fix a primitive $n$-root of unity, the isomorphism $\operatorname{\mathbb{Z}}/n\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times})$ is obtained via the multiplicative structure on $\operatorname{\mathbb{Z}}/n$. We will use a similar approach in the function field case. We first recall the following definition.
Definition 6.2.
(See [20, Page 66-67]) Let $A$ be a finite dimensional commutative $\operatorname{\mathbb{F}}_{q}$-algebra. $A$ is said to be a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra if there exists an $\operatorname{\mathbb{F}}_{q}$-linear functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{q}$ such that the kernel of $\lambda$ contains no nonzero ideal of $A$. We call $\lambda$ a non-degenerate linear functional on $A.$
Example 6.3.
Let $\operatorname{\mathbb{F}}_{q^{r}}$ be a finite extension of $\operatorname{\mathbb{F}}_{q}.$ Then, $\operatorname{\mathbb{F}}_{q^{r}}$ equipped with the canonical trace map $\operatorname{{\rm Tr}}\colon\operatorname{\mathbb{F}}_{q^{r}}\to\operatorname{\mathbb{F}}_{q}$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra.
The following lemma is rather standard.
Lemma 6.4.
Suppose that $A$ is a symmetric $\operatorname{\mathbb{F}}_{q^{r}}$-algebra with a $\operatorname{\mathbb{F}}_{q^{r}}$-linear functional $\lambda:A\to\operatorname{\mathbb{F}}_{q^{r}}.$ Then $A$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra where the linear functional is the composition of $\lambda$ and $\operatorname{{\rm Tr}}\colon\operatorname{\mathbb{F}}_{q^{r}}\to\operatorname{\mathbb{F}}_{q}.$
Proposition 6.5.
Let $A$ be a symmetric finite dimensional $\operatorname{\mathbb{F}}_{q}$-algebra and $\lambda\colon A\to\operatorname{\mathbb{F}}_{q}$ an associated non-degenerate $\operatorname{\mathbb{F}}_{q}$-linear functional. For each $a\in A$, let $\lambda_{a}\colon A\to\operatorname{\mathbb{F}}_{q}$ be the $\operatorname{\mathbb{F}}_{q}$-linear map defined by $\lambda_{a}(b)=\lambda(ab).$ Let $\Phi$ be the map $A\to\operatorname{{\rm Hom}}_{\operatorname{\mathbb{F}}_{q}}(A,\operatorname{\mathbb{F}}_{q})$ sending $a\mapsto\lambda_{a}.$ Then $\Phi$ is an isomorphism.
Proof.
Since $\lambda$ is non-degenerate, $\Phi$ is injective. Furthermore, because $A$ is finite dimensional over $\operatorname{\mathbb{F}}_{q}$, $\dim(A)=\dim(\operatorname{{\rm Hom}}_{\operatorname{\mathbb{F}}_{q}}(A,\operatorname{\mathbb{F}}_{q})).$ Since $\Phi$ is $\operatorname{\mathbb{F}}_{q}$-linear, it must be an isomorphism. ∎
Fix a primitive $p$-root of unity $\zeta_{p}\in\operatorname{\mathbb{C}}^{\times}$. Then, $\operatorname{\mathbb{F}}_{p}$ is (non)-cannonically a subgroup of $\operatorname{\mathbb{C}}^{\times}.$ If $A$ is an $\operatorname{\mathbb{F}}_{q}$-algebra, then $(A,+)$ is a direct sum of several copies of $\operatorname{\mathbb{F}}_{p}$. Therefore
$\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})\cong\operatorname{{\rm Hom}}_{\operatorname{\mathbb{F}}_{p}}(A,\operatorname{\mathbb{F}}_{p}).$
By 6.5 we have the following corollary, which is a direct analog of 6.1.
Corollary 6.6.
Let $A$ be a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra together with a non-degenerate functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{p}.$ For each $a\in A$, let $\lambda_{a}\colon A\to\operatorname{\mathbb{C}}^{\times}$ defined by $\lambda_{a}(b)=\zeta_{p}^{\lambda(ab)}.$ Then $\lambda\in\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times}).$ Furthermore, the map $a\mapsto\lambda_{a}$ gives an isomorphism between $A$ and $\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})$.
We will now focus on the case $A=\operatorname{\mathbb{F}}_{q}[x]/f$. We will show that it is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra (and hence a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra as explained in 6.4). We will show this by constructing an explicit $\operatorname{\mathbb{F}}_{q}$-linear functional on $A.$ We learned about this construction in [19]. Every element $g$ in $\operatorname{\mathbb{F}}_{q}[x]/f$ can be written uniquely in the form
$g=a_{0}(g)+a_{1}(g)+\cdots+a_{n-1}(g)x^{n-1}.$
We define $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ by
$\psi(g)=a_{n-1}(g).$
Proposition 6.7.
Suppose $g\in\operatorname{\mathbb{F}}_{q}[x]/f$ such that $\psi(hg)=0$ for all $h\in\operatorname{\mathbb{F}}_{q}[x]/f.$ Then $g=0$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$
Proof.
Let us write $g=a_{0}(g)+a_{1}(g)+\cdots+a_{n-1}(g)x^{n-1}$. We will prove by induction that $a_{n-k}(g)=0$ for $1\leq k\leq n.$ In fact, since $\psi(g)=0$, we know that $a_{n-1}(g)=0.$ Consequently, the statement is true for $k=1.$ Let us assume that it has been for all $1\leq k\leq m<n$. We claim that it is also true for $m+1$, namely $a_{n-m-1}(g)=0$ as well. In fact, we have
$\begin{split}x^{m-1}g&=x^{m-1}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}+a_{n-m}(g)x^{n}+\cdots+a_{n-1}(g)x^{m+n-2}\\ &=x^{m-1}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}.\end{split}$
Consequently
$a_{n-m-1}(g)=\psi(x^{m-1}g)=0.$
By the induction principle, we conclude that $a_{n-k}=0$ for all $1\leq k\leq n.$
Corollary 6.8.
$\psi$ is a non-degenerate $\operatorname{\mathbb{F}}_{q}$-linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f.$ Consequently, $\operatorname{\mathbb{F}}_{q}[x]/f$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra.
[…]
Proposition 6.9.
There exists a bijection
$\operatorname{\mathbb{F}}_{q}[x]/f\longleftrightarrow\operatorname{{\rm Hom}}((\operatorname{\mathbb{F}}_{q}[x]/f,+),\operatorname{\mathbb{C}}^{\times}),\quad a\longleftrightarrow\{\psi_{a}\},$
where
$\psi_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{C}}^{\times}$
is given by
$\psi_{a}(b)=\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ab))},\forall b\in\operatorname{\mathbb{F}}_{q}[x]/f.$
We remark that over $\operatorname{\mathbb{Z}}$, if $\zeta_{n}$ is a primitive $n$th-root of unity, then for each divisor $m\mid n$, $\zeta_{n}^{\frac{n}{m}}$ is a primitive $m$-root of unity. An analogous statement holds for $\operatorname{\mathbb{F}}_{q}[x]$ as well.
Proposition 6.10.
Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional. Let $g$ be a divisor of $f$ and $\psi_{g}\colon\operatorname{\mathbb{F}}_{q}[x]/g\to\operatorname{\mathbb{F}}_{q}$ be the function defined by
$\psi_{g}(a)=\psi\left(\frac{f}{g}a\right).$
Then $\psi_{g}$ is a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/g.$
Proof.
It is clear from the definition that $\psi_{g}$ is $\operatorname{\mathbb{F}}_{q}$-linear. We only need to show that it is non-degenerate. In fact, suppose to the contrary that the kernel of $\psi_{g}$ contains a non-zero ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ Since $\operatorname{\mathbb{F}}_{q}[x]$ is a PID, $I$ must be of the form $I=\langle h\rangle$ for some $h|g.$ We then see that $\langle h\frac{f}{g}\rangle$ belongs to the kernel of $\psi.$ Because $\psi$ is non-degenerate, this implies that $h\frac{f}{g}=0$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$ In other words, $g|h$ or equivalently $h=0$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ This shows that $I=0$, which is a contradiction. ∎
6.2. Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$
We now introduce the definition of Ramanujan sum over $\operatorname{\mathbb{F}}_{q}[x].$
Definition 6.11.
(Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$) Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be monic polynomials. Let $\psi:\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f$. The Ramanujan sum $c(g,f)$ is defined as
$c(g,f)=c_{\psi}(g,f)=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}.$
Remark 6.12.
We remark also that at first glance, $c(g,f)$ depends on the choice of $\psi.$ However, as we explain in what follows, it does not as long as we make sure that $\psi$ is non-degenerate. This is similar to the case over $\operatorname{\mathbb{Z}}$: Ramanujan sums do not depend on the choice of a primitive $n$th-root of unity. In fact, we will show that there is an explicit formula for $c(g,f)$ similar to the case of Ramanujan sum over $\operatorname{\mathbb{Z}}.$
Remark 6.13.
Ramanujan sums are a special case of Gauss sums as defined and studied in [21, Definition 1]. In fact, they are Gauss sums for the principal Dirichlet characters on $\operatorname{\mathbb{F}}_{q}[x]/f.$ It would be rather interesting if we could define Fekete polynomials for these principal Dirichlet characters (the case over $\operatorname{\mathbb{Z}}$ was studied in [7]).
We first recall the following standard lemma in group theory.
Lemma 6.14.
Let $G$ be a finite group and let $\chi\colon G\to\operatorname{\mathbb{C}}^{\times}$ be a non-trivial character. Then
$\sum_{g\in G}\chi(g)=0.$
Proof.
Since $\chi$ is non-trivial, there exists $h\in G$ such that $\chi(h)\neq 1.$ We then have
$\sum_{g\in G}\chi(g)=\sum_{h\in G}\chi(hg)=\chi(h)\sum_{g\in G}\chi(g).$
Because $\chi(h)\neq 1$, we must have $\sum_{g\in G}\chi(g)=0.$
Proposition 6.15.
Let $\psi$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f$. For each $f\in\operatorname{\mathbb{F}}_{q}[x]$
$c(1,f)=c_{\psi}(1,f)=\mu(f),$
where $\mu$ is the Möbius function on $\operatorname{\mathbb{F}}_{q}[x].$
Proof.
If $f=\prod_{i=1}^{d}f_{i}^{\alpha_{i}}$ where $\alpha_{i}\in\operatorname{\mathbb{N}}$ and $f_{i}$ are irreducible, then by the Chinese remainder theorem and [21, Satz 1] we have
$c_{\psi}(1,f)=\prod_{i=1}^{d}c_{\psi_{i}}(1,f_{i}^{\alpha_{i}}),$
where $\psi_{i}:\operatorname{\mathbb{F}}_{q}[x]/f_{i}^{\alpha_{i}}\to\operatorname{\mathbb{F}}_{q}$ is a non-degenerate linear functional.
Therefore, it is enough to consider the case $f=f_{1}^{a_{1}}$ where $f_{1}$ is irreducible. We have
$c_{\psi}(1,f)$ $=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(a))}=\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(a))}-\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}.$
By 6.14
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(a))}=0.$
Similarly, if $a_{1}\geq 2$ then by 6.14 and 6.10
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=0.$
On the other hand if $a_{1}=1$ then $\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=-1$. We conclude that
$c(1,f)=c_{\psi}(1,f)=c_{\psi}(1,f_{1}^{a_{1}})=\mu(f_{1}^{a_{1}})=\mu(f).\qed$
For general $g$, we have the following theorem.
Theorem 6.16.
Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be monic polynomials. Then (compare with [18, Equation 9])
$c(g,f)=\mu(t)\dfrac{\varphi(f)}{\varphi(t)},\quad\text{where}\quad t=\dfrac{f}{\gcd(f,g)}.$
Proof.
Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional and
$c(g,f)=c_{\psi}(g,f).$
Let $h=\gcd(g,f)$ and $t=\dfrac{f}{h}$. The canonical projection map $(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}\to(\operatorname{\mathbb{F}}_{q}[x]/t)^{\times}$ has kernel of size $\frac{\varphi(f)}{\varphi(t)}.$ Furthermore, for each $a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$, if $a_{1},a_{2}$ are two preimages of $a$ in $(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}$ then $a_{2}=a_{1}+tu$, for some $u\in\operatorname{\mathbb{F}}_{q}[x]/f$. In this case, one has $\psi(ga_{2})=\psi(ag_{1}+gtu)=\psi(ga_{1})+\psi(\frac{gu}{h}f)=\psi(ga_{1})$. Consequently, we have
$c_{\psi}(g,f)$ $=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/t)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}$
$=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(\frac{f}{t}\frac{g}{h}a))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{t}(\frac{g}{h}a))}.$
Here $\psi_{t}:\operatorname{\mathbb{F}}_{q}[x]/t\to\operatorname{\mathbb{F}}_{q}$ is the linear functional defined by $\psi_{t}(b)=\psi(\frac{f}{t}b)=\psi(hb)$. Note that $\psi_{t}$ is non-degenerate.
Since $g/h\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$, we see that
$\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{t}(\frac{g}{h}a))}=\sum_{b\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{t}(t))}=c_{\psi_{t}}(1,t)=\mu(t).$
Here the last equality follows from 6.15. We conclude that
$c(g,f)=c_{\psi}(g,f)=\frac{\varphi(f)}{\varphi(t)}\mu(t).\qed$
6.3. Spectrum of $G_{f}(D)$
Let $G$ be an abelian group and $S$ a symmetric subset of $G$ such that $0\not\in S$. The adjacency matrix of the Cayley graph $\Gamma(G,S)$ is a $G$-circulant matrix. By the $G$-circulant theorem, the spectrum of $\Gamma(G,S)$ is the collection of the sums $\sum_{s\in S}\chi(s)$, where $\chi$ runs over the set of all characters of $G$ (see [14, Section 1.2]). When $G=\operatorname{\mathbb{F}}_{q}[x]/f$, we know by 6.9 that the character of $G$ is parameterized by $G$ itself. Consequently, we have the following proposition.
Proposition 6.17.
Let $S$ be a subset of $\operatorname{\mathbb{F}}_{q}[x].$ Then the spectrum of the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ is given by the set
$\left\{\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/f}.$
Let us now focus on the case of gcd-graphs; i.e., $S=S_{D}$. We have the following lemma.
Lemma 6.18.
Let $h\mid f$ be a monic divisor of $f.$ Then for each $g\in\operatorname{\mathbb{F}}_{q}[x]$
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f,\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=c\left(g,\frac{f}{h}\right).$
Proof.
For $a\in\operatorname{\mathbb{F}}_{q}[x]$, $\gcd(a,f)=h$ if and only if $a=hb$ where $\gcd(b,\frac{f}{h})=1.$ Therefore, the above sum can be rewritten as
$\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(a,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(bgh))}=\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(b,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}.$
Here $\psi_{h}:\operatorname{\mathbb{F}}_{q}[x]/(f/h)\to\operatorname{\mathbb{F}}_{q}$ is the functional given by $\psi_{h}(x)=\psi(hx).$ By 6.10, $\psi_{h}$ is a non-generate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f/g).$ Therefore, we conclude that
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f,\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(b,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}=c\left(g,\frac{f}{h}\right).$
By 6.17 and 6.18, we have the following theorem.
Theorem 6.19.
Let $f$ be a monic polynomial and $D=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}\mid f.$ Then, the spectrum of $S_{f}(D)$ is given by the set
$\left\{\sum_{i=1}^{k}c\left(g,\frac{f}{f_{i}}\right)\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/f}.$
Corollary 6.20.
All eigenvalues of the gcd-graph $G_{f}(D)$ are integers.
As we discussed in the introduction, it is known that a $\operatorname{\mathbb{Z}}/n$-circulant graph has an integral spectrum if and only if it is a gcd-graph (see [29, Theorem 7.1] ). One may ask whether the same statement is true for graphs associated with $\operatorname{\mathbb{F}}_{q}[x]/f.$ The answer is no in general. In fact, we have the following general observation.
Proposition 6.21.
Let $S$ be a symmetric subset of $\operatorname{\mathbb{F}}_{q}[x]/f$ such that $0\not\in S.$ Suppose further that $\operatorname{\mathbb{F}}_{p}^{\times}S=S.$ Then, the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ has an integral spectrum.
Proof.
We define the following equivalence relation on $\operatorname{\mathbb{F}}_{q}[x]/f$. For $u,v\in\operatorname{\mathbb{F}}_{q}[x]/f$, we say that $u\sim v$ if $u=kv$ where $k\in\operatorname{\mathbb{F}}_{p}^{\times}.$ By 6.17, it is enough to show that for each $g\in\operatorname{\mathbb{F}}_{q}[x]/f$, $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$. Because $\operatorname{\mathbb{F}}_{p}^{\times}S=S$, this sum can be written as
$\sum_{[s]\in S/\sim}\left(\sum_{k\in\operatorname{\mathbb{F}}_{p}^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gks))}\right)=\sum_{[s]\in S/\sim}\left(\sum_{k\in\operatorname{\mathbb{F}}_{p}^{\times}}(\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))})^{k}\right).$
We know that
$\sum_{k\in\operatorname{\mathbb{F}}_{p}^{\times}}(\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))})^{k}=\begin{cases}-1&\text{if }\operatorname{{\rm Tr}}(\psi(gs))\neq 0\\ p-1&\text{if }\operatorname{{\rm Tr}}(\psi(gs))=0.\end{cases}$
This shows that $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$ for each $g\in\operatorname{\mathbb{F}}_{q}[x]/f.$
Remark 6.22.
Let $\operatorname{\mathbb{F}}_{q}$ be a finite field such that $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{p}.$ Let $f=x$. In this case $\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{q}.$ Let $V$ be proper $\operatorname{\mathbb{F}}_{p}$-subspace of $\operatorname{\mathbb{F}}_{q}$ and $S=V\setminus\{0\}.$ Then by 6.21, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ has an integral spectrum even though it is not a gcd-graph.
We wonder whether the converse of 6.21 holds (perhaps, under some mild assumptions).
7. Perfect gcd-graphs
A graph $G$ is said to be perfect if, for every induced subgraph $H$ of $G$, the chromatic number of $H$ equals the size of its maximum clique. Perfect graphs play a fundamental role in the study of graph coloring and cliques. They encompass several important families of graphs and provide a unified framework for results relating to colorings and cliques within these families. Moreover, many central problems in combinatorics can be rephrased as questions about whether certain associated graphs are perfect (see [9, 10, 24]). For these reasons, it seems interesting to study whether $G_{f}(D)$ is perfect.
In [1, Theorem 9.5], the authors classify all perfect unitary Cayley graphs associated with finite commutative rings. Specifically, for a ring $R=R_{1}\times R_{2}\times\cdots\times R_{t}$ where $R_{i}$ are finite local rings such that $|R_{1}|\leq|R_{2}|\leq\ldots\leq|R_{t}|$, the unitary Cayley graph on $R$ is perfect if and only if one of the following conditions hold
  1. (1)
    The residue field of $R_{1}$ is $2$. In this case, $G_{R}$ is bipartite and hence perfect.
  2. (2)
    $R$ is either a local ring or a product of two local rings; i.e, $t\leq 2.$
A direct consequence of [1, Theorem 9.5] for the unitary Cayley graph over $\operatorname{\mathbb{Z}}$ is the following.
Corollary 7.1.
The unitary Cayley graph on $\operatorname{\mathbb{Z}}/N$ is perfect if and only if one of the following conditions holds.
  1. (1)
    $2\mid N$.
  2. (2)
    $\omega(N)\leq 2$ where $\omega(N)$ is the number of distinct irreducible factors of $N.$
Over the polynomial ring $\operatorname{\mathbb{F}}_{q}[x]$, we have an analogous statement.
Corollary 7.2.
The unitary Cayley graph $G_{f}(\{1\})$ on $\operatorname{\mathbb{F}}_{q}[x]/f$ is perfect if and only if one of the following conditions holds.
  1. (1)
    $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $gcd(f,x(x+1))\neq 1.$
  2. (2)
    $\omega(f)\leq 2$ where $\omega(f)$ is the number of distinct irreducible factors of $f.$
We will use 7.2 together with patterns of $G_{f}(D)$ discovered through our experiments in Sagemath and the Python package Networkx in order to find some sufficient conditions for $G_{f}(D)$ to be non-perfect. In particular, we will exploit the fact that certain induced subgraphs of $G_{f}(D)$ are naturally isomorphic to the unitary Cayley graphs on some quotient rings of $R=\operatorname{\mathbb{F}}_{q}[x]/f.$ More precisely, by 5.4, we have the following observation.
Proposition 7.3.
Suppose $g\in D$ such that $g\nmid f_{i}$ for all $i$ such that $f_{i}\neq g.$ Then the induced graph on $I_{g}$ is naturally isomorphic to the unitary Cayley graph $G_{f/g}(\{1\}).$ Furthermore, if $\omega(f/g)\geq 3$ and $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$, then $G_{f/g}(D)$ is not perfect and hence $G_{f}(D)$ is not perfect.
We discuss a case where we can apply 7.3 rather directly.
Proposition 7.4.
Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and ${D}=\{f_{1},f_{2},\ldots,f_{k}\}$ be a subset of divisors of $f$. Suppose that the following conditions hold.
  1. (1)
    $\deg(f_{i})\geq 1.$
  2. (2)
    $f_{i}$’s are pairwisely relatively prime; i.e, $\gcd(f_{i},f_{j})=1$ for all $i\neq j.$
  3. (3)
    $k\geq 3.$
  4. (4)
    $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}.$
Then $G_{f}(D)$ is not a perfect graph.
Proof.
If $k\geq 4$ then
$\omega(f/f_{1})\geq\omega(f_{2})+\omega(f_{3})+\omega(f_{4})\geq 3.$
By applying 7.3 for $g=f_{1}$, we conclude that $G_{f}(D)$ is not perfect. Let us assume now that $k=3$. By the same argument, it is enough to consider the case $\omega(f_{1})=\omega(f_{2})=\omega(f_{3})=1.$ Furthermore, if $f\neq f_{1}f_{2}f_{3}$, then there exists an index $i\in\{1,2,3\}$ such that $\omega(f/f_{i})\geq 3.$ As a result, $G_{f}(D)$ is not perfect by 7.3. Let us now consider the case $f=f_{1}f_{2}f_{3}$. In this case
$\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{q}[x]/f_{1}\times\operatorname{\mathbb{F}}_{q}[x]/f_{2}\times\operatorname{\mathbb{F}}_{q}[x]/f_{3}.$
Under this isomorphism, we can identify $V(G_{f}(D))$ as the set of all triples $(a_{1},a_{2},a_{3})$ where $a_{i}\in\operatorname{\mathbb{F}}_{q}[x]/f_{i}$ for $1\leq i\leq 3$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if there exists an index $i\in\{1,2,3\}$ such that $a_{i}=b_{i}$ and $(a_{j}-b_{i})\in(\operatorname{\mathbb{F}}_{q}[x]/f_{j})^{\times}$ if $j\neq i$. Using Sagemath, we can find the following induced $5$-cycle in $G_{f}(D)$
$(0,0,0)\to(\alpha,\alpha,0)\to(1,\alpha,1)\to(1,1,\alpha)\to(\alpha,1,0),$
where $\alpha\in\operatorname{\mathbb{F}}_{q}\setminus\{0,1\}.$ This show that $G_{f}(D)$ is not perfect. ∎
The case $k=2$ is a bit more challenging. After some experiments with Sagemath, we found the following statement.
Proposition 7.5.
Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and ${D}=\{f_{1},f_{2}\}$ be a subset of divisors of $f$ such that $\gcd(f_{1},f_{2})=1$. Suppose that the following conditions hold.
  1. (1)
    $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$.
  2. (2)
    $\omega(f_{1})\geq 2$, $\omega(f_{2})\geq 1$.
Then, $G_{f}(D)$ is not perfect.
Proof.
Let us assume to the contrary that $G_{f}(D)$ is perfect. Since $\gcd(f_{1},f_{2})=1$, we have $f_{1}f_{2}\mid f$ and hence
$\omega(f)\geq\omega(f_{1})+\omega(f_{2}).$
If $\omega(f)>\omega(f_{1})+\omega(f_{2})$ then $\omega(f/f_{2})>\omega(f_{1})\geq 2.$ By 7.3, we know $G_{f}(D)$ is not perfect. Therefore, we must have $\omega(f)=\omega(f_{1})+\omega(f_{2}).$
Let $g=f_{1}f_{2}.$ By 5.6, the ideal generated by $g$ is a homogeneous set in $G_{f}(D)$ and furthermore $G_{f}(D)$ is isomorphic to the wreath product $G_{g}(\{f_{1},f_{2}\})*G_{f/g}(\emptyset).$ Since $G_{f}(D)$ is perfect, $G_{g}(\{f_{1},f_{2}\})$ is perfect as well. Because $\omega(f_{1})\geq 2$, we can find $h_{1},h_{2}\in\operatorname{\mathbb{F}}_{q}[x]$ such that $f_{1}=h_{1}h_{2}$ and $\gcd(h_{1},h_{2})=1.$ By the Chinese remainder theorem
$\operatorname{\mathbb{F}}_{q}[x]/g\cong\operatorname{\mathbb{F}}_{q}[x]/h_{1}\times\operatorname{\mathbb{F}}_{q}[x]/h_{2}\times\operatorname{\mathbb{F}}_{q}[x]/f_{2}.$
Under this isomorphism, we can identify $\operatorname{\mathbb{F}}_{q}[x]/g$ with the set of tuples $(a_{1},a_{2},a_{3})$ such that $a_{1}\in\operatorname{\mathbb{F}}_{q}[x]/h_{1},a_{2}\in\operatorname{\mathbb{F}}_{q}[x]/h_{2},a_{3}\in\operatorname{\mathbb{F}}_{q}[x]/f_{2}.$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if one of the following conditions happens
  1. (1)
    $a_{1}=b_{1},a_{2}=b_{2}$ and $a_{3}-b_{3}\in(\operatorname{\mathbb{F}}_{q}[x]/f_{2})^{\times}$,
  2. (2)
    $a_{3}=b_{3}$ and $a_{i}-b_{i}\in(\operatorname{\mathbb{F}}_{q}[x]/h_{i})^{\times}$ for $i\in\{1,2\}.$
Using Sagemath, we can find the following $7$-cycle in $G_{g}(\{f_{1},f_{2}\})$
$(0,0,0)\to(1,1,0)\to(\alpha,0,0)\to(\alpha,0,1)\to(0,1,1)\to(\alpha,\alpha,1)\to(0,0,1),$
where $\alpha\in\operatorname{\mathbb{F}}_{q}\setminus\{0,1\}.$ We conclude that $G_{g}(\{f_{1},f_{2}\})$ is not a perfect graph, which is a contradiction. ∎
Remark 7.6.
Curious readers might wonder why we choose a $7$-cycle in the proof of 7.5 instead of choosing a $5$-cycle as in the proof of 7.4. The reason is that when $\omega(f_{1})=2,\omega(f_{2})=1$ and $f=f_{1}f_{2}$, our code cannot find $5$-cycle in $G_{f}(D).$ It seems interesting to investigate whether this is always the case.
Remark 7.7.
The only remaining case that we miss is when $\omega(f_{1})=\omega(f_{2})=1.$ We distinguish the following cases.
  1. (1)
    If $f=f_{1}f_{2}$ where both $f_{1},f_{2}$ are irreducible then $G_{f}(D)$ is isomorphic to the complement of $G_{f}(\{1\})$ which is perfect by 7.2. Therefore, $G_{f}(D)$ is perfect as well.
  2. (2)
    On the other hand, if either $f_{1}$ or $f_{2}$ is reducible, our code can always find a $7$-cycle in $G_{f}(D).$ Unfortunately, we cannot find a universal pattern in this case. It would be quite interesting to solve this puzzle completely.
8. Induced subgraphs of gcd-graphs
A theorem of Erdős and Evans (see [11]) says that every graph $G$ is an induced subgraph of the unitary Cayley graph on $\operatorname{\mathbb{Z}}/n$ for some squarefree $n.$ Using this result, it is shown in [16] that for a given finite graph $G$ and a finite field $F$, $G$ is an induced subgraph of the unitary graph of a matrix algebra $M_{d}(F)$ for some value of $d$ (they also provide some precise upper-bound on $d$ when $G$ is the complete graph $K_{m}$). In light of these results, it seems interesting to ask whether a graph $G$ can be realized as an induced subgraph of $G_{f}(D)$ for some choice of $f,D$ and $\operatorname{\mathbb{F}}_{q}.$ Under some rather mild conditions, the answer is yes as we will show below. First, we introduce the following observation.
Lemma 8.1.
If $G$ is an induced subgraph of $H$, then $G$ is also an induced subgraph of the product $H\times K_{|G|}$ where $K_{|G|}$ is the complete graph on $|G|$-nodes.
Proof.
Let $f\colon G\to H$ be a graph morphism that makes $G$ into an induced subgraph of $H.$ Let us index $V(G)=\{v_{1},v_{2},\ldots,v_{|G|}\}$. Let $\hat{f}\colon G\to H\times K_{|G|}$ be the map defined by
$\hat{f}(v_{i})=(f(v_{i}),i).$
We can see that $\hat{f}$ is a graph morphism which turns $G$ into an induced subgraph of $H\times K_{|G|}.$
We are now ready to prove an analog of Erdős-Evans’s theorem in the function fields case.
Proposition 8.2.
Let $G$ be a fixed graph and $q$ a fixed prime power. There exists a positive integer $r$ such that for each $d\geq r$, we can find $F_{d}\in\operatorname{\mathbb{F}}_{q}[x]$ satisfying the following conditions
  1. (1)
    $\omega(F_{d})=d$,
  2. (2)
    $G$ is an induced subgraph of the unitary Cayley graph $G_{F_{d}}(\{1\})$.
Proof.
By Erdős and Evans’s theorem [11], there exists a squarefree integer $n\in\operatorname{\mathbb{N}}$ such that $G$ is an induced subgraph of the unitary graph $G_{n}$ on $\operatorname{\mathbb{Z}}/n.$ Let $r=\omega(n)$; i.e, $n=p_{1}p_{2}\ldots p_{r}$ be the prime factorization of $n$. Then
$G_{n}\cong\prod_{i=1}^{r}G_{p_{i}}\cong\prod_{i=1}^{r}K_{p_{i}}.$
For each $i\in\operatorname{\mathbb{N}}$ and $m_{i}\in\operatorname{\mathbb{N}}$, there exists a polynomial $h_{i}$ of degree $m_{i}$ such that $h_{i}$ is irreducible over $\operatorname{\mathbb{F}}_{q}[x].$ We then see that $G_{h_{i}}(\{1\})$ is isomorphic to the complete graph $K_{q^{m_{i}}}.$ Let us choose $m_{i}$ so that $q^{m_{i}}>n$ for all $i$. For each $d\geq r$, let $F_{d}=h_{1}h_{2}\ldots h_{d}.$ We then have
$G_{F_{d}}(\{1\})\cong\prod_{i=1}^{d}G_{h_{i}}(\{1\})\cong\prod_{i=1}^{d}K_{q^{m_{i}}}.$
Since $q^{m_{i}}>n$, we know that $G_{n}$ is an induced subgraph of $G_{F_{r}}(\{1\}).$ By 8.1 , $G$ is an induced subgraph of $G_{F_{d}}(\{1\})$ as well.
Corollary 8.3.
Let $G$ be a fixed graph and $k$ a fixed positive integer. Then, there exist a polynomial $f$ and a subset $D=\{f_{1},f_{2},\ldots,f_{k}\}$ of divisors of $f$ such that $G$ is an induced subgraph of $G_{f}(D).$
Proof.
Let $h_{0}$ be an arbitrary polynomial in $\operatorname{\mathbb{F}}_{q}[x].$ By 8.2, there exists a positive integer $d\geq k$ and a polynomial $h$ of the form $h=h_{1}h_{2}\ldots h_{d}$ such that $G$ is an induced subgraph of $G_{h}(\{1\}).$ Let us choose $f=h_{0}h$ and
$\{f_{1},f_{2},\ldots,f_{k}\}=\{h_{1},h_{2},\ldots,h_{k}\}.$
Let $I$ be the ideal generated by $h_{0}$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$ By 5.4 the induced graph on $I$ is naturally isomorphic to the unitary Cayley graph $G_{h}(\{1\})$. This shows that $G$ is an induced subgraph of $G_{f}(D).$
One may wonder whether the following stronger form of 8.3 holds. Let $G$ be a fixed graph and $f_{1},f_{2},\ldots,f_{k}$ fixed polynomials. Does there exist a polynomial $f$ such that
  1. (1)
    $f_{i}\mid f$ for all $1\leq i\leq k$,
  2. (2)
    $G$ is an induced subgraph of $G_{f}(D)$ where $D=\{f_{1},f_{2},\ldots,f_{k}\}$.
In general, the answer is no. There seem to be some subtle constraints. We discuss here a particle one. Let $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})$ and assume further that $g\neq f_{i}$ for all $1\leq i\leq k$. If such $f$ exists, then $g\mid f$ and there is a canonical map
$\Phi:\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g.$
If we take a subset $S\subset\operatorname{\mathbb{F}}_{q}[x]/f$ such that $|S|>|\operatorname{\mathbb{F}}_{q}[x]/g|$, then there exists $a,b\in S$ such that $a\neq b$ and $\Phi(a)=\Phi(b).$ Consequently, $\Phi(a-b)=0$ or $g\mid a-b.$ By definition, $(a,b)\not\in E(G_{f}(D)).$ Consequently, we have the following upper bound for the clique number of $G_{f}(D)$
$\omega(G_{f}(D))\leq|\operatorname{\mathbb{F}}_{q}[x]/g|.$
This shows that if $\omega(G)>|\operatorname{\mathbb{F}}_{q}[x]/g|$ then $G$ cannot be an induced subgraph of $G_{f}(D).$
We can overcome the constraint by enlarging the base field $\operatorname{\mathbb{F}}_{q}.$
Proposition 8.4.
Let $G$ be a fixed graph. Let $f_{1},f_{2},\ldots,f_{k}$ be fixed polynomials in $\operatorname{\mathbb{F}}_{q}[x]$. Then, there exist a finite extension $\operatorname{\mathbb{F}}_{q^{m}}$ of $\operatorname{\mathbb{F}}_{q}$ and a polynomial $f\in\operatorname{\mathbb{F}}_{q^{m}}[x]$ such that
  1. (1)
    $f_{i}\mid f$ for all $1\leq i\leq k$,
  2. (2)
    $G$ is an induced subgraph of $G_{f}(D)$ where $D=\{f_{1},f_{2},\ldots,f_{k}\}$ and we consider $f$ as an element of $\operatorname{\mathbb{F}}_{q^{m}}[x].$
Proof.
By Erdős and Evans’s theorem, there exists a squarefree number $n$ such that $G$ is an induced subgraph of the unitary Cayley graph $G_{n}.$ Suppose that $r=\omega(n)$ and $n=p_{1}p_{2}\ldots p_{r}$ be the prime factorization of $n.$ Then, as explained in the proof of 8.2, $G_{n}\cong\prod_{i=1}^{r}K_{p_{i}}.$ Our goal is to show that we can find $\operatorname{\mathbb{F}}_{q^{m}}$ and $f$ such that $G_{n}$ is an induced subgraph of $G_{f}(D)$ where we consider $f$ as an element in $\operatorname{\mathbb{F}}_{q^{m}}[x].$
Let $m$ be a positive integer such that $q^{m}>n$ and $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})\in\operatorname{\mathbb{F}}_{q}[x].$ By Galois theory, there exists a polynomial $h\in\operatorname{\mathbb{F}}_{q}[x]$ with at least $r$ distinct irreducible factors and $\gcd(h,g)=1.$ Let $f=hg$. Let $I$ be the ideal in $\operatorname{\mathbb{F}}_{q^{m}}[x]/f$ generated by $f_{1}.$ Then, by 5.4, the induced subgraph on $I$ is isomorphic to the unitary Cayley graph $G_{\bar{f}}(\{1\})$ where $\bar{f}=f/f_{1}.$ Suppose that $\bar{f}=g_{1}^{a_{1}}g_{2}^{a_{2}}\ldots g_{t}^{a_{t}}$ be the factorization of $f/f_{1}$ over $\operatorname{\mathbb{F}}_{q^{m}}[x].$ Then, by the choice of $f$, $t\geq r.$ By the Chinese remainder theorem, we know that $\operatorname{\mathbb{F}}_{q^{m}}^{t}$ is a subring of $\operatorname{\mathbb{F}}_{q^{m}}[x]/\bar{f}.$ Therefore, $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}=\prod_{i=1}^{t}K_{q^{m}}$ is an induced subgraph of $G_{\bar{f}}(\{1\})$. Since $t\geq r$ and $q^{m}\geq n$, by 8.1 $G_{n}$ is an induced subgraph of $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}$. Consequently, $G_{n}$ is an induced subgraph of $G_{f}(D)$ as well. ∎
Acknowledgements
We thank Professor Torsten Sander for the enlightening correspondence about gcd-graphs and related arithmetic graphs. We are grateful to Professor Emanuel Kolwaski for some helpful comments on Fekete polynomials and for writing the book [19]. We thank Professor Ki-Bong Nam for his interest in this topic and his feedback on the first draft. Finally, we thank Minh-Tam Trinh for his discussions, encouragement, and support at the initial stage of this work.

### v2

@@
-On the gcd graphs over polynomial rings and related topics
+On the gcd-graphs over polynomial rings
 Abstract.
-Gcd-graphs over the ring of integers modulo $n$ are a natural generalization of unitary Cayley graphs.
-The study of these graphs has foundations in various mathematical fields, including number theory, ring theory, and representation theory.
-Using the theory of Ramanujan sums, it is known that these gcd-graphs have integral spectra; i.e., all their eigenvalues are integers.
-In this work, inspired by the analogy between number fields and function fields, we define and study gcd-graphs over polynomial rings with coefficients in finite fields.
-We establish some fundamental properties of these graphs, emphasizing their analogy to their counterparts over $\operatorname{\mathbb{Z}}.$
+[…]
 Key words and phrases:
-Cayley graphs, Function fields, Ramanujan sums, Symmetric algebras, Integral graphs, Prime graphs, Perfect graphs.
+Gcd-graphs, Cayley graphs, prime graphs, perfect graphs, connected graphs, bipartite graphs, Ramanujan sums, integral spectra
 2020 Mathematics Subject Classification:
@@
 Let $n$ be a positive integer.
-The unitary Cayley graph on the ring of integers modulo $n$ is defined as the Cayley graph $G_{n}={\rm Cay}(\operatorname{\mathbb{Z}}/n,U_{n})$ where $\operatorname{\mathbb{Z}}/n$ is the ring of integers modulo $n$ and $U_{n}=(\operatorname{\mathbb{Z}}/n)^{\times}$ its unit group.
-More specifically, $G_{n}$ is equipped with the following data.
+The unitary Cayley graph on the ring of integers modulo $n$ is defined as the Cayley graph $G_{n}={\rm Cay}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},U_{n})$ where $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ is the ring of integers modulo $n$ and $U_{n}=(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}})^{\times}$ its unit group.
+More specifically, $G_{n}$ is equipped with the following data:
   1. (1)
-    The vertex set of $G_{n}$ is $\operatorname{\mathbb{Z}}/n$.
+    The vertex set of $G_{n}$ is $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$;
   2. (2)
@@
 For example, [1] studies the unitary Cayley graph of a finite commutative ring.
-[15] generalizes this study further to finite rings which are not necessarily commutative.
-In [26], a subset of the authors studies various arithmetic and graph-theoretic properties of the $p$-unitary Cayley graph as defined by [27, 28] (when $p=1$, these graphs are precisely unitary Cayley graphs).
+In [16], the authors generalize this study further to finite rings which are not necessarily commutative.
+In [27], a subset of the authors studies various arithmetic and graph-theoretic properties of a subclass of $k$-unitary Cayley graphs as defined in [29, 31] (when $k=1$, these graphs are precisely unitary Cayley graphs).
 We refer the interested readers to [4, 9, 19, 27] and the references therein for some further topics in this line of research.
-As explained in [18, Section 4], a particularly intriguing arithmetical property of $G_{n}$ is that its spectrum can be described by the theory of circulant matrices and Ramanujan sums.
+As explained in [19, Section 4], a particularly intriguing arithmetical property of $G_{n}$ is that its spectrum can be described by the theory of circulant matrices and Ramanujan sums.
 A consequence of this fact is that the unitary Cayley graph has an integral spectrum; i.e., all of its eigenvalues are integers.
@@
   1. (1)
-    The vertex set of $G_{n}(D)$ is $\operatorname{\mathbb{Z}}/n.$
+    The vertex set of $G_{n}(D)$ is $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}.$
   2. (2)
     Two vertices $a,b\in G_{n}$ are adjacent if and only if $\gcd(a-b,n)\in D.$
-We remark that the unitary Cayley graph $G_{n}$ is nothing but $G_{n}(\{1\}).$ It is known that the spectrum of $G_{n}(D)$ is a summation of various Ramanujan sums (see [18, Section 4]).
+We remark that the unitary Cayley graph $G_{n}$ is nothing but $G_{n}(\{1\}).$ We also note that the gcd-graph $G_{n}(D)$ is loopless since $D$ does not contain $n$.
+Additionally, the graph $G_{n}(D)$ is undirected since $\gcd(a-b,n)=\gcd(b-a,n)$.
+It is known that the spectrum of $G_{n}(D)$ is a summation of various Ramanujan sums (see [19, Section 4]).
 In particular, all of its eigenvalues are integers.
 It turns out that, the converse is also true.
-Theorem 1.1.
-(See [29, Theorem 7.1]) A $\operatorname{\mathbb{Z}}/n$-circulant graph $G$ is an integral graph if and only if $G=G_{n}(D)$ for some $D.$
-Let $\operatorname{\mathbb{F}}_{q}$ be a finite field.
-As observed by Andre Weil in [30], there is a strong analogy between the ring $\operatorname{\mathbb{Z}}$ of integers and the ring $\operatorname{\mathbb{F}}_{q}[x]$ of polynomials with coefficients in $\operatorname{\mathbb{F}}_{q}$ (and more generally, between number fields and function fields).
-Consequently, it is of reasonable interest to define and investigate gcd-graphs over $\mathbb{F}_{q}[x]$.
-Fortunately, the analogous definition of gcd-graphs over $\mathbb{F}_{q}[x]$ is relatively straightforward.
-More specifically, let $f\in\operatorname{\mathbb{F}}_{q}[x]$ be a non-zero element in $\operatorname{\mathbb{F}}_{q}[x]$.
-Let $D=\{f_{1},f_{2},\ldots,f_{k}\}$ be a subset of the set of divisors $\operatorname{{\rm Div}}(f)$ of $f$; i.e, $f_{i}\mid f$ for all $1\leq i\leq k.$ Let
-$S_{D}=\{g\in\operatorname{\mathbb{F}}_{q}[x]/f|\gcd(g,f)\in D\}.$
-Definition 1.2.
-We denote by $G_{f}(D)$ the gcd-Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S_{D}).$ More precisely, it is the graph equipped with the following data
-  1. (1)
-    The vertex set of $G_{f}(D)$ is the finite ring $\operatorname{\mathbb{F}}_{q}[x]/f.$
-  2. (2)
-    Two vertices $u,v$ are adjacent if and only if $u-v\in S_{D}$. In other words, $\gcd(u-v,f)\in D.$
+Theorem 1.1 (See [32, Theorem 7.1]).
+An $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$-circulant graph $G$ is an integral graph if and only if $G=G_{n}(D)$ for some $D.$
+[…]
+$S_{D}=\{g\in\operatorname{\mathbb{F}}_{q}[x]/(f):\gcd(g,f)\in D\}.$
 Remark 1.2.
@@
 In our case, however, we can make a canonical choice for the great common divisor by requiring it to be a monic polynomial.
-Therefore, unless we explicitly state, we will assume throughout this article that all involved polynomials are monic.
+Therefore, unless we explicitly state the contrary, we will assume throughout this article that all involved polynomials are monic.
+Definition 1.3.
+The gcd-graph $G_{f}(D)$ is the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S_{D}).$ More precisely, it is the graph equipped with the following data:
+  1. (1)
+    The vertex set of $G_{f}(D)$ is the finite ring $\operatorname{\mathbb{F}}_{q}[x]/(f)$,
+  2. (2)
+    Two vertices $u,v$ are adjacent if and only if $u-v\in S_{D}$. In other words, $\gcd(u-v,f)\in D$}.
+[…]
 Example 1.4.
-Fig. 1 shows the graph $G_{f}(D)$ where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$ and $D=\{x,x+1\}$. It is a regular graph of degree $4$
-Figure 1. The gcd-Cayley graph $G_{x(x+1)}(\{x,x+1\})$
-1.1. Outline
-In this article, we study some foundational properties of gcd-graphs defined over $\mathbb{F}_{q}[x]$.
+Fig. 1 shows the graph $G_{f}(D)$ where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$ and $D=\{x,x+1\}$}. We have $S_{D}=\{x,2x,x+1,2(x+1)\}$. The graph $G_{f}(D)$ is a regular graph of degree $4.$
+Figure 1. The gcd-graph $G_{x(x+1)}(\{x,x+1\})$
+Outline
+In this article we introduce and study some foundational properties of gcd-graphs defined over $\mathbb{F}_{q}[x]$.
 In Section 2, we recall some standard definitions in graph theory that will be used throughout the article.
 Section 3 investigates the question of when a gcd-graph is connected and anti-connected.
-In Section 4, we provide the necessary and sufficient conditions for a gcd-graph to be bipartite.
-Section 5 explores another graph-theoretic property of gcd-graphs: their prime-property.
-Although we are not able to provide a complete answer, we present some rather general conditions under which this property holds.
+(Recall that a graph is anti-connected if its complement is connected.) In Section 4, we provide the necessary and sufficient conditions for a gcd-graph to be bipartite.
+We note that our investigation provides additional insights into the properties of Cayley graphs.
+For example, in recent works [29, 30], the authors examine the connectedness and bipartiteness of a related class of graphs known as generalized Paley graphs.
+Section 5 explores another graph-theoretic property of gcd-graphs: their primeness property.
+Here, we describe the necessary and sufficient conditions for a gcd-graph to be prime.
+We then apply this result to give an explicit criterion for a gcd-graph to be prime under some mild assumptions on the set $D$.
 Section 6 examines the spectra of gcd-graphs using symmetric algebras and Ramanujan sums.
@@
 As a by-product, we prove an analogous result to a theorem of Erdős and Evans in the number field case.
-1.2. Code
+Code
 The code that we develop to generate gcd-graphs and do experiments on them can be found at [24].
@@
 Definition 2.1 (The complete graph $K_{n}$).
-$K_{n}$ is the graph on $n$ vertices which are pairwisely adjacent.
+The complete graph $K_{n}$ is the graph on $n$ vertices which are pairwisely adjacent.
+We now review two types of graph products that appear in our work, starting with the tensor product.
 Definition 2.2 (Tensor product of graphs).
@@
     Two vertices $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G\times H$ if and only if $(g,g^{\prime})\in E(G)$ and $(h,h^{\prime})\in E(H).$
-Definition 2.3 (Wreath product).
-Let $G,H$ be two graphs. We define the wreath product of $G$ and $H$ as the graph $G*H$ with the following data
+Next, we recall the definition of the wreath product of graphs.
+Definition 2.3 (Wreath product of graphs).
+Let $G,H$ be two graphs. We define the wreath product of $G$ and $H$ as the graph GHG*H with the following data:
   1. (1)
-    The vertex set of $G*H$ is the Cartesian product $V(G)\times V(H)$,
+    The vertex set of GHG*H is the Cartesian product $V(G)\times V(H)$,
   2. (2)
-    $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in $G*H$ if either $(g,g^{\prime})\in E(G)$ or $g=g^{\prime}$ and $(h,h^{\prime})\in E(H)$.
-Definition 2.4 (Graph morphism).
+    $(g,h)$ and $(g^{\prime},h^{\prime})$ are adjacent in GHG*H if either $(g,g^{\prime})\in E(G)$ or $g=g^{\prime}$ and $(h,h^{\prime})\in E(H)$.
+Definition 2.4 (Graph morphisms).
 Let $G_{1}$ and $G_{2}$ be two graphs.
-We define a graph morphism between $G_{1}$ and $G_{2}$ to be a map from $V(G_{1})$ to $V(G_{2})$ $f$ that preserves edges.
+We define a graph morphism between $G_{1}$ and $G_{2}$ to be a map $f$ from $V(G_{1})$ to $V(G_{2})$ that preserves edges.
 More precisely, if $u,v\in V(G_{1})$ are adjacent in $G_{1}$, then $f(u),f(v)$ are adjacent in $G_{2}$.
-Definition 2.5 (Induced subgraph).
-Let $G_{1}$ and $G_{2}$ be two graphs. We say that $G_{1}$ is an induced subgraph of $G_{2}$ if there exists a graph morphism $f:G_{1}\to G_{2}$ such that the following conditions hold.
+Finally, we recall the definition of induced subgraphs.
+Definition 2.5 (Induced subgraphs).
+Let $G_{1}$ and $G_{2}$ be two graphs. We say that $G_{1}$ is an induced subgraph of $G_{2}$ if there exists a graph morphism $f\colon G_{1}\to G_{2}$ such that the following conditions hold.
   1. (1)
@@
   2. (2)
-    For $u,v\in G_{1}$, $u,v$ are adjacent if and only if $f(u)$ and $f(v)$ are adjacent in $G_{2}.$
-3. Connectedness of gcd graphs
+    For $\in G_{1}, u,vu,v$ are adjacent if and only if $f(u)$ and $f(v)$ are adjacent in $G_{2}.$
+3. Connectedness of gcd-graphs
 In this section, we will give the necessary and sufficient conditions for $G_{f}(D)$ to be connected.
 Unlike the number field case where these conditions are rather straightforward, the function field case is a bit more complicated.
-This is partially due to the fact that the unitary Cayley graph on $\operatorname{\mathbb{F}}_{q}[x]/f$ is not always connected.
+This is partially due to the fact that the unitary Cayley graph on $\operatorname{\mathbb{F}}_{q}[x]/(f)$ is not always connected.
 We start our discussion with the following simple observation.
@@
 Let $f_{0}=\gcd(f_{1},\ldots,f_{k})$.
-We observe that if two vertices $u$ and $v$ are adjacent then $f_{0}$ divides $u-v$.
-Now since $G_{f}(D)$ is connected, there is a path connecting 00 and $1$.
-This implies that $f_{0}$ divides $1-0=1$.
+We observe that if two vertices $u$ and $v$ are adjacent then $f_{0}$ divides uvu-v.
+Now since $G_{f}(D)$ is connected, there is a path connecting 0 and $1$.
+This implies that $f_{0}$ divides 10=11-0=1.
 Hence $\gcd(f_{1},\ldots,f_{k})=1$.
 ∎
-In the case of $\operatorname{\mathbb{Z}}$, the converse of 3.1 is true as well; i.e.
-the gcd-graph $G_{n}(D)$, where $D=\{d_{1},d_{2},\ldots,d_{k}\}$ is connected if and only if ${\rm gcd}(d_{1},d_{2},\ldots,d_{k})=1.$ In the case of $\operatorname{\mathbb{F}}_{q}[x]$, this condition is not sufficient.
-For example, let $D=\{1\}$, $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2},f=x(x+1).$ Then $\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{2}\times\operatorname{\mathbb{F}}_{2}.$ In this case
+[…]
 $G_{f}(\{1\})\cong K_{2}\times K_{2},$
@@
 Proof.
+Let $R=\operatorname{\mathbb{F}}_{q}[x]/(f)$.
 We need to show that $R=\langle S_{D}\rangle$ the abelian group generated by $S_{D}$.
-Let $a\in R=\operatorname{\mathbb{F}}_{q}[x]/f.$ Because $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, we can find $a_{1},a_{2},\ldots,a_{k}\in R$ such that
+Let $a\in R.$ Because $\gcd(f_{1},f_{2},\ldots,f_{k})=1$, we can find $a_{1},a_{2},\ldots,a_{k}\in R$ such that
 $a=\sum_{i=1}^{k}a_{i}f_{i}.$
-Since $G_{f}(\{1\})$ is connected, for each $1\leq i\leq d$, we can find write
+Since $G_{f}(\{1\})$ is connected, for each $1\leq i\leq k$, we can find write
 $a_{i}=\sum_{j=1}^{n_{i}}m_{ij}s_{ij},$
@@
 $a=\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}m_{ij}s_{ij}f_{i}.$
-By definition, $\gcd(s_{ij}f_{i},f)=f_{i}\in D$. This shows that $a\in\langle S_{D}\rangle$. Since this is true for all $a$, we conclude that $R=\langle S_{D}\rangle$. ∎
+Since $f_{i}\mid f$, we see that $\gcd(s_{ij}f_{i},f)=f_{i}\in D$. This shows that $a\in\langle S_{D}\rangle$. Since this is true for all $a$, we conclude that $R=\langle S_{D}\rangle$. ∎
 We can classify all $f$ such that $G_{f}(\{1\})$ is not connected.
-Proposition 3.3.
-(See [8, Lemma 4.33]) $G_{f}(\{1\})$ is not connected if and only if $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$
+Proposition 3.3 (See [9, Lemma 4.33]).
+$G_{f}(\{1\})$ is not connected if and only if $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$
 By 3.2 and 3.3, we have the following corollary.
@@
 Proof.
-Let $m=\gcd(h,f).$ We need to show that $\gcd(h,g)=\gcd(m,g).$ We first claim that $\gcd(h,g)\mid\gcd(m,g).$ By definition, $\gcd(h,g)\mid g.$ We also have $\gcd(h,g)\mid\gcd(h,f)=m.$ Therefore, $\gcd(h,g)\mid\gcd(m,g).$
+Let $m=\gcd(h,f).$ We need to show that $\gcd(h,g)=\gcd(m,g).$ We first claim that $\gcd(h,g)\mid\gcd(m,g).$ By definition, $\gcd(h,g)\mid g.$ We also have $\gcd(h,g)\mid\gcd(h,f)=m$}.
+Therefore, $\gcd(h,g)\mid\gcd(m,g).$
 Conversely, we claim that $\gcd(m,g)\mid\gcd(h,g).$ This is clear since $m\mid h$.
@@
 Proposition 3.6.
-Let $g$ be a divisor of $f$ and $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g$ be the canonical projection map. Let
-$\overline{D}=\{\gcd(f_{i},g)|1\leq i\leq k\}.$
+Let $g$ be a divisor of $f$ and $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/g$ be the canonical projection map. Let
+$\overline{D}=\{\gcd(f_{i},g):1\leq i\leq k\}.$
 Then $\Phi_{f,g}(S_{D})\subseteq S_{\overline{D}}$. Consequently, $\Phi_{f,g}\colon G_{f}(D)\to G_{g}(\overline{D})$ is a graph morphism.
@@
 Proof.
-Let $\bar{a},\bar{b}$ be two vertices in $G_{g}(\overline{D}).$ Since $\Phi_{f,g}$ is surjective, we can find $a,b\in\operatorname{\mathbb{F}}_{q}[x]/f$ such that $\Phi_{f,g}(a)=\bar{a},\Phi_{f,g}(b)=\bar{b}.$ Since $G_{f}(D)$ is connected, there is a path $P$ from $a$ to $b.$ By 3.6, $\Phi_{f,g}(P)$ is a path from $\bar{a}$ to $\bar{b}.$ This shows that $G_{g}(\overline{D})$ is connected as well.
+Let $\bar{a},\bar{b}$ be two vertices in $G_{g}(\overline{D}).$ Since $\Phi_{f,g}$ is surjective, we can find $a,b\in\operatorname{\mathbb{F}}_{q}[x]/(f)$} such that $\Phi_{f,g}(a)=\bar{a},\Phi_{f,g}(b)=\bar{b}.$ Since $G_{f}(D)$ is connected, there is a path $P$ from $a$ to $b.$ By 3.6, $\Phi_{f,g}(P)$ is a path from $\bar{a}$ to $\bar{b}.$ This shows that $G_{g}(\overline{D})$ is connected as well.
 ∎
@@
 Theorem 3.8.
-Suppose that $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$ Let $\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/(x(x+1))$ be the canonical projection map and $\bar{D}$ as described above.
-Then $G_{f}(D)$ is connected if and only if
+Suppose that $\operatorname{\mathbb{F}}_{q}=\operatorname{\mathbb{F}}_{2}$ and $x(x+1)\mid f.$ Let
+$\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(x(x+1))$
+be the canonical projection map and $\bar{D}$ as described above. Then $G_{f}(D)$ is connected if and only if
   1. (1)
@@
   2. (2)
-    the graph $G_{x(x+1)}(\bar{D})$ is connected. This condition is equivalent to $|\overline{D}\setminus\{x(x+1)\}|\geq 2.$
+    the graph $G_{x(x+1)}(\bar{D})$ is connected. This condition is equivalent to D ∖{x(x+1) }2.
 Proof.
 By 3.7, we know that if $G_{f}(D)$ is connected then $(1)$ and $(2)$ holds.
-Conversley, let us assume that $(1)$ and $(2)$ both hold.
+Conversely, let us assume that $(1)$ and $(2)$ both hold.
 We will show that $G_{f}(D)$ is connected.
-This is equivalent to showing that $R:=\operatorname{\mathbb{F}}_{q}[x]/f=\langle S_{D}\rangle.$
+This is equivalent to showing that $R:=\operatorname{\mathbb{F}}_{q}[x]/(f)=\langle S_{D}\rangle.$
 The key idea of this proof is similar to the proof for 3.2. The main difficulty is to deal with the fact that $G_{f}(\{1\})$ is not connected in this case. We will do this step by step.
-We first claim that if $g\in R$ such that $x(x+1)\mid g$ then $g=s_{1}+s_{2}$ where $s_{1},s_{2}\in R^{\times}.$ For an element $s\in R$, $s$ is a unit if and only $\Phi^{\text{ss}}(s)\in(R^{\text{ss}})^{\times}$ where $R^{\text{ss}}=R/J(R)$ is the semisimplification of $R$ (see [8, Proposition 4.30]).
+We first claim that if $g\in R$ such that $x(x+1)\mid g$ then $g=s_{1}+s_{2}$ where $s_{1},s_{2}\in R^{\times}.$ For an element $\in R, ss$ is a unit if and only $\Phi^{\text{ss}}(s)\in(R^{\text{ss}})^{\times}$ where $R^{\text{ss}}=R/J(R)$ is the semisimplification of $R$ (see [9, Proposition 4.30]).
 Therefore, for this statement, we can assume that $R=R^{\text{ss}}$; namely $f$ is a squarefree polynomial.
-If we write $f=x(x+1)f_{1}$ where $\gcd(x(x+1),f_{1})=1$ then we have the isomorphism
-$R\cong\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))\times\operatorname{\mathbb{F}}_{2}[x]/f_{1}.$
-Under this isomorphism, $g$ is sent to $(0,g_{1})$ where $g_{1}\in\operatorname{\mathbb{F}}_{2}[x]/f_{1}.$ Since $F_{2}[x]/f_{1}$ is a product of fields of order bigger than $2$, every element in it can be written as the sum of two units; say $g_{1}=t_{1}+t_{2}$ where $t_{1},t_{2}\in(\operatorname{\mathbb{F}}_{2}[x]/f)^{\times}.$ We then see that $g=s_{1}+s_{2}$ where $s_{1}=(1,t_{1})$ and $s_{2}=(1,t_{2})$.
+If we write $f=x(x+1)h$ where $\gcd(x(x+1),h)=1$ then we have the isomorphism
+$R\cong\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))\times\operatorname{\mathbb{F}}_{2}[x]/(h).$
+Under this isomorphism, $g$ is sent to $(0,g_{1})$ where $g_{1}\in\operatorname{\mathbb{F}}_{2}[x]/(h).$ Since $F_{2}[x]/(h)$ is a product of fields of order bigger than $2$, every element in it can be written as the sum of two units; say $g_{1}=t_{1}+t_{2}$ where $t_{1},t_{2}\in(\operatorname{\mathbb{F}}_{2}[x]/(f))^{\times}.$ We then see that $g=s_{1}+s_{2}$ where $s_{1}=(1,t_{1})$ and $s_{2}=(1,t_{2})$.
 By definition $s_{1},s_{2}\in R^{\times}.$
@@
 (3.1) $g=\sum_{i=1}^{k}(s_{1i}f_{i}+s_{2i}f_{i}).$
-Since $s_{1},s_{2}\in R^{\times}$, $s_{1i}f_{i},s_{2i}f_{i}\in S_{D}.$ This shows that $g\in\langle S_{D}\rangle.$
+Since $s_{1i},s_{2i}\in R^{\times}$, $s_{1i}f_{i},s_{2i}f_{i}\in S_{D}.$ This shows that $g\in\langle S_{D}\rangle.$
 Finally, let $g$ now be an arbitrary element in $R$.
-We claim that $g\in\langle S_{D}\rangle.$ By our assumption, the graph $G_{x(x+1)}(S_{\bar{D}})$ is connected, 00 and $\Phi_{f,x(x+1)}(g)$ are connected by a path.
+We claim that $g\in\langle S_{D}\rangle.$ By our assumption, the graph $G_{x(x+1)}(S_{\bar{D}})$ is connected, 0 and $\Phi_{f,x(x+1)}(g)$ are connected by a path.
 Consequently, we can write
@@
 Remark 3.9.
-The proof of 3.8 also implies that for the case $\operatorname{\mathbb{F}}_{2}[x]$ and $x(x+1)\mid f$, the map $\Phi_{f,g}:\operatorname{\mathbb{F}}_{2}[x]/f\to\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))$ has the property that $a$ and $b$ belong the same connected component in $G_{f}(D)$ if and only $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ belong to the same connected component in $G_{x(x+1)}(\overline{D}).$ In fact, by 3.7, if $a$ and $b$ are connected by a path then $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path.
+The proof of 3.8 also implies that for the case $\operatorname{\mathbb{F}}_{2}[x]$ and $x(x+1)\mid f$}, the map $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{2}[x]/(f)\to\operatorname{\mathbb{F}}_{2}[x]/(x(x+1))$ has the property that $a$ and $b$ belong to the same connected component in $G_{f}(D)$ if and only $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ belong to the same connected component in $G_{x(x+1)}(\overline{D}).$ In fact, by 3.7, if $a$ and $b$ are connected by a path then $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path.
 Conversely, if $\Phi_{f,x(x+1)}(a)$ and $\Phi_{f,x(x+1)}(b)$ are connected by a path in $G_{x(x+1)}(\overline{D})$, we can write
@@
 where $n_{i}\in\operatorname{\mathbb{Z}}$ and $\gcd(h_{1},f)\in D$.
-Because $x(x+1)R\subseteq\langle S_{D}\rangle$, we conclude that that $a-b\in\langle S_{D}\rangle.$ By definition, $a$ and $b$ are connected by a path in $G_{f}(D)$.
-We conclude that $G_{f}(D)$ and $G_{x(x+1)}(\overline{D})$ has the same number of connected component. In particular, the number of connected component in $G_{f}(D)$ is at most $2$
+From the proof of 3.8, we know that $x(x+1)R\subseteq\langle S_{D}\rangle$.
+Therefore, we conclude that $a-b\in\langle S_{D}\rangle.$ By definition, $a$ and $b$ are connected by a path in $G_{f}(D)$.
+We conclude that $G_{f}(D)$ and $G_{x(x+1)}(\overline{D})$ has the same number of connected component. In particular, the number of connected components in $G_{f}(D)$ is at most $2.$
+Example 3.10.
+Fig.
+2 shows the graph $G_{f}(D)$ where $f=x^{2}(x+1)\in\operatorname{\mathbb{F}}_{2}[x]$ and $D=\{x,x+1\}$}.
+We have $S_{D}=\{x,x+1,x^{2}+1\}$.
+The graph $G_{f}(D)$ is a regular graph of degree $3.$ Furthermore, by 3.8, it is connected.
+Figure 2. The gcd-Cayley graph $G_{x^{2}(x+1)}(\{x,x+1\})$
 Remark 3.11.
@@
 A graph $G=(V,E)$ is called a bipartite graph if $V$ can be partitioned into two disjoint sets $V_{1},V_{2}$ such that for every edge $(u,v)\in E$ either $u\in V_{1},v\in V_{2}$ or $u\in V_{2},v\in V_{1}$.
+In this case, we write $V=V_{1}\sqcup V_{2}$ and we call $V_{1},V_{2}$ independent sets in $G.$
 Bipartite graphs model various real-life situations such as job assignments, resource allocation, stable marriage matching, gene-disease association, and much more.
@@
 We refer the reader to [3] for some further discussions on this topic.
-The goal of this section is to classify all $G_{f}(D)$ which is bipartite.
-We remark that while our focus is on gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x]$, everything we discuss in this section applies to gcd-graphs over $\operatorname{\mathbb{Z}}$ as well (see 4.5).
-To the best of our knowledge, even over $\operatorname{\mathbb{Z}}$, this topic has not been explored in the literature.
-Therefore, in addition to studying this problem for its own merits, we also hope to address a gap in the literature.
+[…]
 We start with the following lemma.
@@
     Let
-    $I=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 0\pmod{2}\right\},$
-    $I_{1}=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 1\pmod{2}\right\},$
+    $I=\left\{\sum_{i}n_{i}h_{i}:n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 0\pmod{2}\right\}$,
+    $I_{1}=\left\{\sum_{i}n_{i}h_{i}:n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}^{k}n_{i}\equiv 1\pmod{2}\right\}$.
     Then $I$ is an subgroup of index $2$ in $(R,+).$ Furthermore, $I$ and $I_{1}$ are independent set such that $V(G_{f}(D))=I\bigsqcup I_{1}.$
@@
 Proof.
-  1. (1)
-    Let us consider the first part of this lemma. Because $G_{f}(D)$ is bipartite, we can write
+(1) Let us consider the first part of this lemma. Because $G_{f}(D)$ is bipartite, we can write
 $V(G_{f}(D))=A\bigsqcup B$
@@
 We claim that $A=I$.
-By the proof of [4, Proposition 2.6], we know that $A$ is a subgroup of $(R,+)$ of index $2$ By definition, if $h_{i}\in S_{D}$, then $(0,h_{i})\in E(G_{f}(D))$.
+By the proof of [5, Proposition 2.6], we know that $A$ is a subgroup of $(R,+)$ of index $2.$ By definition, if $h_{i}\in S_{D}$, then $(0,h_{i})\in E(G_{f}(D))$.
 Because $A$ and $B$ are disjoint independent set in $G_{f}(D)$, $h_{i}\in B$.
 Furthermore, since $A$ is a subgroup of index $2$ in $R$, $h_{i}+h_{j}\in A$ for all $h_{i},h_{j}\in S_{D}.$ Consequently, we must have $I\subseteq A.$
-    By our assumption that $G_{f}(D)$ is connected we know that $R=\langle S_{D}\rangle$ and hence $I\sqcup I_{1}=\langle S_{D}\rangle=R.$ Clearly if $a\in R\setminus I=I_{1}$ then $2a\in I$.
+By our assumption that $G_{f}(D)$ is connected we know that $R=\langle S_{D}\rangle$ and hence $I\sqcup I_{1}=\langle S_{D}\rangle=R$}.
+Clearly if $a\in R\setminus I=I_{1}$ then $2a\in I$.
 From this, we can see that $I$ is a subgroup whose index is at most $2$ in $R$.
 Additionally, because $I\subseteq A$ and $A$ has index $2$ in $R$, we must have that $I=A$ and $I_{1}=B.$
-  2. (2)
-    Suppose that $G_{f}(\{1\})$ is connected.
+(2) Suppose that $G_{f}(\{1\})$ is connected.
 We will show that $I$ is an ideal in $R$ as well.
@@
 Specifically, we note that for each $s\in R^{\times}$, $sI=I.$ Since $\langle R^{\times}\rangle=R$, this shows that $aI\subseteq I$ for all $a\in R.$ We conclude that $I$ is an ideal in $R$.
-  3. (3)
-    Finally, part $(3)$ follows from Equation Eq. 3.1.
+(3) Finally, part $(3)$ follows from Equation Eq. 3.1. ∎
 Corollary 4.2.
@@
 Proof.
-If $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$ or $\gcd(x(x+1),f)=1$, we know that $G_{f}(\{1\})$ is connected by 3.3.
-Furthermore, $\operatorname{\mathbb{F}}_{q}[x]/f$ has no ideal of index $2$ By 4.1, we conclude that $G_{f}(D)$ is not bipartite.
+If $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{2}$ or $\gcd(x(x+1),f)=1$, we know that $G_{f}(\{1\})$ is connected by $3.$ Furthermore, $\operatorname{\mathbb{F}}_{q}[x]/(f)$ has no ideal of index $2.$ By 4.1, we conclude that $G_{f}(D)$ is not bipartite.
 ∎
@@
 Theorem 4.3.
-Suppose that $f$ is a polynomial in $\operatorname{\mathbb{F}}_{2}[x]$ such that $\gcd(x(x+1),f)\not\in\{1,x(x+1)\}.$ Then $G_{f}(D)$ is bipartite if and only if $\gcd(x(x+1),f)\nmid f_{i}$ for all $1\leq i\leq k.$
+Suppose that $f$ is a polynomial in $\operatorname{\mathbb{F}}_{2}[x]$ such that $\gcd(x(x+1),f)\not\in\{1,x(x+1)\}$}.
+Then $G_{f}(D)$ is bipartite if and only if $\gcd(x(x+1),f)\nmid f_{i}$ for all $1\leq i\leq k.$
 Proof.
@@
 In other words, $x\nmid f_{i}$ for all $1\leq i\leq k.$
-Finally, let us consider the trickest case where $x(x+1)\mid f.$
+Finally, let us consider the trickiest case where $x(x+1)\mid f.$
 Theorem 4.4.
 Suppose that $x(x+1)\mid f$ and that $G_{f}(D)$ is connected. Let
-$\overline{D}=\{\gcd(f_{i},x(x+1))|f_{i}\in D.\}$
+$\overline{D}=\{\gcd(f_{i},x(x+1)):f_{i}\in D\}.$
 Then $G_{f}(D)$ is bipartite if and only if $|\overline{D}|=2.$
@@
 and
-$B=\{g\in R|g(0)\neq g(1)\}=x+A.$
+$B=\{g\in R:g(0)\neq g(1)\}=x+A.$
 We can check that if $a_{1},a_{2}\in A$ then $\gcd(a_{1}-a_{2},x(x+1))\in\{1,x(x+1)\}.$ By definition, $(a_{1},a_{2})\not\in E(G_{f}(D)).$ This shows that $A$ is an independent set in $G_{f}(D)$.
 Similarly, $B$ is an independent set in $G_{f}(D)$ as well.
-We conclude that $G_{f}(D)$ is not bipartite.
+We conclude that $G_{f}(D)$ is bipartite.
 Conversely, suppose $G_{f}(D)$ is bipartite. We claim that $|\overline{D}|=2.$ By 4.1, the subgroup
-$I=\left\{\sum_{i}n_{i}h_{i}|n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}n_{i}\equiv 0\pmod{2}\right\},$
-is an independent set in $G_{f}(D).$ Furthermore, by part $(3)$ of 3.7, we know that $x(x+1)R\subset I.$ Because 00 is not connected to any node in $I$, we must have $x(x+1)R\cap S_{D}=\emptyset$.
+$I=\left\{\sum_{i}n_{i}h_{i}:n_{i}\in\operatorname{\mathbb{Z}},h_{i}\in S_{D}\text{ and}\quad\sum_{i=1}n_{i}\equiv 0\pmod{2}\right\}$,
+is an independent set in $G_{f}(D).$ Furthermore, by part $(3)$ of 3.7, we know that $x(x+1)R\subset I.$ Because 0 is not connected to any node in $I$, we must have $x(x+1)R\cap S_{D}=\emptyset$}.
 Consequently $x(x+1)\not\in\overline{D}.$ Suppose to the contrary that $|\overline{D}|\geq 3$.
-We then have $\overline{D}=\{1,x,x+1\}.$ This implies that $G_{x(x+1)}(\overline{D})$ is the complete graph $K_{4}.$ In $K_{4}$, there is a path of length $2$ from 00 to any other vertices.
+We then have $\overline{D}=\{1,x,x+1\}.$ This implies that $G_{x(x+1)}(\overline{D})$ is the complete graph $K_{4}.$ In $K_{4}$, there is a path of length $2$ from 0 to any other vertices.
 From this property, we conclude that for each $g\in R$, we can write
@@
 We then conclude that $g-h_{1}-h_{2}\in x(x+1)R\subset I$.
-Since $h_{1}+h_{2}\in I$, we conclude that $g\in I$ as well.
+By the definition of $I$, $h_{1}+h_{2}\in I$, we conclude that $g\in I$ as well.
 This shows that $I=R$, which is a contradiction.
@@
 Remark 4.5.
-By an almost identical argument, we can show that the gcd graph $G_{n}(D)$ over $\operatorname{\mathbb{Z}}$ with $D=\{d_{1},d_{2},\ldots,d_{k}\}$ and $\gcd(d_{1},d_{2},\ldots,d_{k})=1$, is bipartite if and only if $2\mid n$ and $2\nmid d_{i}$ for all $1\leq i\leq k.$
-5. Primeness of gcd graphs
+[…]
+5. Primeness of gcd-graphs
 For a given graph $G$, a homogeneous set in $G$ is a set $X$ of vertices of $G$ such that every vertex in $V(G)\setminus X$ is adjacent to either all or none of the vertices in $X$.
@@
 In [9], in collaboration with some other graph theorists, we completely classify prime unitary Cayley graphs on finite commutative rings.
-In this section, we study the problem for the gcd graph $G_{f}(D).$ We remark that for a graph $G$, a connected component of $G$ (or of its complement) is a homogeneous set in $G$.
+In this section, we study the problem for the gcd-graph $G_{f}(D).$ We remark that for a graph $G$, a connected component of $G$ (or of its complement) is a homogeneous set in $G$.
 Therefore, while studying the primeness of $G$, it is safe to assume that $G$ is connected and anti-connected.
@@
 Proposition 5.1.
-Let $a\in(\operatorname{\mathbb{F}}_{q}[t]/f)^{\times}.$ Let $m_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/f$ be the multiplication by $a$ map.
+Let $a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}.$ Let $m_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(f)$ be map induced by the multiplication by $a$.
 Then $m_{a}$ induces an automorphism on $G_{f}(D).$
 Proof.
-Since $a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}$, $m_{a}$ is an automorphism of $(\operatorname{\mathbb{F}}_{q}[x]/f,+)$.
+Since $a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}$, $m_{a}$ is an automorphism of $(\operatorname{\mathbb{F}}_{q}[x]/(f),+)$.
 Furthermore, $m_{a}$ preserves $S_{D}$; i.e $aS_{D}=S_{D}.$ As a result, $m_{a}$ is an automorphism of $G_{f}(D).$
@@
   2. (2)
-    There exists a non-trivial ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/f$ such that $I$ is a homogeneous set in $G_{f}(D).$
+    There exists a non-trivial ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $I$ is a homogeneous set in $G_{f}(D).$
 Proof.
@@
 Let us now prove $(1)$ implies $(2)$.
-The proof that we discuss here is quite similar to the one that we gave for [8, Theorem 4.1] and 3.2.
-Because $G_{f}(D)$ is not prime, we can find a maximal homogenous set $H$ containing $0$ By [8, Theorem 3.4], we know that $H$ is a subgroup of $\operatorname{\mathbb{F}}_{q}[x]/f$.
+The proof that we discuss here is quite similar to the one that we gave for [9, Theorem 4.1] and 3.2.
+Because $G_{f}(D)$ is not prime, we can find a maximal homogenous set $H$ containing $0.$ By [9, Theorem 3.4], we know that $H$ is a subgroup of $\operatorname{\mathbb{F}}_{q}[x]/(f)$.
 We claim that it is an ideal as well.
-By 5.1, we conclude that $aH=H$ for all $a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}.$ Furthermore, since $G_{f}(\{1\})$ is connected, $aH\subseteq H$ for all $a\in\operatorname{\mathbb{F}}_{q}[x]/f$ as well.
-This shows that $H$ is an ideal in $\operatorname{\mathbb{F}}_{q}[x]/f.$
+By 5.1, we conclude that $aH=H$ for all $a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}.$ Furthermore, since $G_{f}(\{1\})$ is connected, $aH\subseteq H$ for all $a\in\operatorname{\mathbb{F}}_{q}[x]/(f)$ as well.
+This shows that $H$ is an ideal in $\operatorname{\mathbb{F}}_{q}[x]/(f).$
 By 5.2, in order to study the prime property of $G_{f}(D)$, it is essentially equivalent to classify $g$ such that the ideal generated by $g$ is a homogenous set in $G_{f}(D).$ In order to do so, we first introduce the following lemmas.
@@
 We remark that the relation $a-bt_{1}=a_{1}c$ and the fact that $\gcd(a,b)=1$ imply that $\gcd(a_{1},b)=1.$ By the Chinese remainder theorem, we can find $t_{2}$ such that $\gcd(a_{1}-bt_{2},\frac{f}{c})=1$}.
+∎
 Lemma 5.4.
-Let $g$ be a divisor of $f$. Let $I_{g}$ be the ideal in $R=\operatorname{\mathbb{F}}_{q}[x]/f$ generated by $g.$ The induced graph on $I_{g}$ is isomorphic to $G_{f/g}(D_{g})$ where
-$D_{g}=\left\{\frac{f_{i}}{g}|f_{i}\in D,g\mid f_{i}\right\}.$
+Let $g$ be a divisor of $f$. Let $I_{g}$ be the ideal in $R=\operatorname{\mathbb{F}}_{q}[x]/(f)$ generated by $g.$ The induced graph on $I_{g}$ is isomorphic to $G_{f/g}(D_{g})$ where
+$D_{g}=\left\{\frac{f_{i}}{g}:f_{i}\in D,g\mid f_{i}\right\}.$
 Proof.
@@
 Therefore we have a natural map $I_{g}\to\operatorname{\mathbb{F}}_{q}[x]/(f/g)$ sending $gm\mapsto m.$ Furthermore, for two elements $ga,gb\in I_{g}$, we have
-$\gcd(ga-gb,f)=g\gcd(a-b,\frac{f}{g}).$
+$\gcd(ga-gb,f)=g\gcd(a-b,\tfrac{f}{g}).$
 Therefore, we see that $\gcd(ga-gb,f)\in D$ if and only $\gcd(a-b,\frac{f}{g})\in D_{g}.$ From this, we conclude that the induced graph on $I_{g}$ is naturally isomorphic to $G_{f/g}(D_{g})$. ∎
 Theorem 5.5.
-Let $g\mid f$ be a divisor of $f$ and $I$ the ideal in $\operatorname{\mathbb{F}}_{q}[x]/f$ generated by $g.$ Let
-$D_{1}=\left\{f_{i}\in D\mid g\nmid f_{i}\right\},\quad D_{2}=\left\{f_{i}\in D|g\mid f_{i}\right\}.$
+Let $g\mid f$ be a divisor of $f$ and $I$ the ideal in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ generated by $g.$ Let
+$D_{1}=\left\{f_{i}\in D:g\nmid f_{i}\right\},\quad D_{2}=\left\{f_{i}\in D:g\mid f_{i}\right\}.$
 As in 3.6 let
-$\overline{D_{1}}=\left\{\gcd(f_{i},g)\mid f_{i}\in D_{1}\right\},\quad\widetilde{D_{2}}=\left\{\frac{f_{i}}{g}\mid f_{i}\in D_{2}\right\}.$
+$\overline{D_{1}}=\left\{\gcd(f_{i},g):f_{i}\in D_{1}\right\},\quad\widetilde{D_{2}}=\left\{\frac{f_{i}}{g}:f_{i}\in D_{2}\right\}.$
 Then, the following statements are equivalent.
   1. (1)
-    $I$ is a homogeneous set in $G_{f}(D)$.
+    II is a homogeneous set in $G_{f}(D)$.
   2. (2)
-    $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$ where $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g$ is the canonical projection map.
-Furthermore, if one of the above equivalent conditions holds, $G_{f}(D)$ is isomorphic to the wreath product $G_{g}(D_{1})*G_{f/g}(\widetilde{D_{2}}).$
+    $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$ where $\Phi_{f,g}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(g)$ is the canonical projection map.
+[…]
 Proof.
@@
 In other words, suppose that $g$ satisfies the condition that $\Phi_{f,g}^{-1}(\overline{D_{1}})\cap\operatorname{{\rm Div}}(f)=D_{1}$.
-We claim that the ideal $I$ generated by $g$ is a homogeneous set in $G_{f}(D).$ In fact, let $a\not\in I$ be an arbitrary element in $\operatorname{\mathbb{F}}_{q}[x]/f$ and suppose that $a$ is adjacent to an element in $I$.
-By a translation, we can assume that $a$ is adjacent to 00 in $G_{f}(D).$ We claim that $a$ is adjacent to all elements in $I$ as well.
+We claim that the ideal $I$ generated by $g$ is a homogeneous set in $G_{f}(D).$ In fact, let $a\not\in I$ be an arbitrary element in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ and suppose that $a$ is adjacent to an element in $I$.
+By a translation, we can assume that $a$ is adjacent to 0 in $G_{f}(D).$ We claim that $a$ is adjacent to all elements in $I$ as well.
 Let $gt$ be an element in $I$.
@@
 Since $(a,0)\in E(G_{f}(D))$ we know that $\gcd(a,f)\in D.$ Because $a\not\in I$, we know further that $\gcd(a,f)\in D_{1}$.
-Additionally, by 3.5, we must have $\gcd(a,g)=\gcd(\gcd(a,f),g)\in\overline{D_{1}}.$ Again, by 3.5, we have
+Additionally, by 3.5, we must have $\gcd(a,g)=\gcd(\gcd(a,f),g)\in\overline{D_{1}}$}.
+Again, by 3.5, we have
 $\gcd(\gcd(a-gt,f),g)=\gcd(a-gt,g)=\gcd(a,g)\in\overline{D_{1}}.$
@@
 Suppose this is not the case.
-We can find a non-constant irreducible polynomial $h$ of such that $\gcd(h,\frac{g}{f_{i}})=1$ and $h|\frac{f}{f_{i}}.$ We then see that $\gcd(hf_{i},g)=f_{i}=\gcd(f_{i},g).$ This implies that $hf_{i}\in D$.
+We can find a non-constant irreducible polynomial $h$ of such that $\gcd(h,\frac{g}{f_{i}})=1$ and $h\mid\frac{f}{f_{i}}.$ We then see that $\gcd(hf_{i},g)=f_{i}=\gcd(f_{i},g).$ This implies that $hf_{i}\in D$.
 Since $f_{i}\mid hf_{i}$ and $f_{i}\neq hf_{i}$, this leads to a contradiction.
@@
 The proof of 5.6 relies crucially on the divisibility relationship between $f_{i}$ and $f_{j}.$ It seems important to study this relationship systematically.
-We are studying this problem in a work in progress (see [22]).
+Since the completion of this paper, we have made some partial progress on this problem (see [23]).
 Remark 5.9.
-While we primarily focus on gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x]$ in this section, most statements have a straightforward analog for gcd-graphs over $\operatorname{\mathbb{Z}}.$
-6. Spectrum of gcd graphs
-The spectrum of gcd-graphs over $\operatorname{\mathbb{Z}}$ is described by the theory of Ramanujan sums (see [18, Section 4]), which in turn is a special case of Gauss sums (see [21]).
-As explained in [7], these sums are precisely values of Fekete polynomials at certain $n$-roots of unity.
-One might wonder whether such a similar statement for the spectrum of gcd-graphs holds in the context of function fields.
-While we are not able to find an analog of Fekete polynomials over function fields, the theory of Ramanujan sums does have an interesting analogy as we will explain in this section.
-This, however, is sufficient for us to describe explicitly the spectrum of gcd-graphs over $\operatorname{\mathbb{F}}_{q}[x].$
+[…]
+6. Spectrum of gcd-graphs
+[…]
 6.1. Symmetric algebras
-A key point in the theory of $\operatorname{\mathbb{Z}}/n$-circulant graph is the fact that the character group of $\operatorname{\mathbb{Z}}/n$ is isomorphic to $\operatorname{\mathbb{Z}}/n$:
-$\operatorname{\mathbb{Z}}/n\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times}).$
-This isomorphism can be obtained as follows.
-Fix a primitive $n$-root of unity $\zeta_{n}$ in $\operatorname{\mathbb{C}}$.
-Let $\chi_{1}\colon\operatorname{\mathbb{Z}}/n\to\operatorname{\mathbb{C}}^{\times}$ be the character defined by $\chi_{1}(m)=\zeta_{n}^{m}$ for all $m\in\operatorname{\mathbb{Z}}/n.$ For each $a\in\operatorname{\mathbb{Z}}/n$, let $\chi_{a}=\chi_{1}^{a}$ be the character of $\operatorname{\mathbb{Z}}/n$ defined by $\chi_{a}(b)=\zeta_{n}^{ab}=\chi_{1}^{a}(b).$ The following proposition is standard.
+A key point in the theory of $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$-circulant graph is the fact that the character group of $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ is isomorphic to $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$:
+$\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},\operatorname{\mathbb{C}}^{\times}).$
+[…]
 Proposition 6.1.
-The map $a\mapsto\chi_{1}^{a}$ gives an isomorphism between $\operatorname{\mathbb{Z}}/n$ and $\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times}).$
-In summary, once we fix a primitive $n$-root of unity, the isomorphism $\operatorname{\mathbb{Z}}/n\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n,\operatorname{\mathbb{C}}^{\times})$ is obtained via the multiplicative structure on $\operatorname{\mathbb{Z}}/n$.
+The map $a\mapsto\chi_{1}^{a}$ gives an isomorphism between $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ and $\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},\operatorname{\mathbb{C}}^{\times}).$
+In summary, once we fix a primitive $n$-root of unity, the isomorphism $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}\cong\operatorname{{\rm Hom}}(\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}},\operatorname{\mathbb{C}}^{\times})$} is obtained via the multiplicative structure on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$.
 We will use a similar approach in the function field case.
@@
 Definition 6.2.
-(See [20, Page 66-67]) Let $A$ be a finite dimensional commutative $\operatorname{\mathbb{F}}_{q}$-algebra.
+(See [21, Page 66-67]) Let $A$ be a finite dimensional commutative $\operatorname{\mathbb{F}}_{q}$-algebra.
 $A$ is said to be a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra if there exists an $\operatorname{\mathbb{F}}_{q}$-linear functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{q}$ such that the kernel of $\lambda$ contains no nonzero ideal of $A$.
@@
 ∎
-Fix a primitive $p$-root of unity $\zeta_{p}\in\operatorname{\mathbb{C}}^{\times}$.
-Then, $\operatorname{\mathbb{F}}_{p}$ is (non)-cannonically a subgroup of $\operatorname{\mathbb{C}}^{\times}.$ If $A$ is an $\operatorname{\mathbb{F}}_{q}$-algebra, then $(A,+)$ is a direct sum of several copies of $\operatorname{\mathbb{F}}_{p}$.
-Therefore
+[…]
 $\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})\cong\operatorname{{\rm Hom}}_{\operatorname{\mathbb{F}}_{p}}(A,\operatorname{\mathbb{F}}_{p}).$
@@
 Corollary 6.6.
-Let $A$ be a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra together with a non-degenerate functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{p}.$ For each $a\in A$, let $\lambda_{a}\colon A\to\operatorname{\mathbb{C}}^{\times}$ defined by $\lambda_{a}(b)=\zeta_{p}^{\lambda(ab)}.$ Then $\lambda\in\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times}).$ Furthermore, the map $a\mapsto\lambda_{a}$ gives an isomorphism between $A$ and $\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})$.
-We will now focus on the case $A=\operatorname{\mathbb{F}}_{q}[x]/f$.
-We will show that it is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra (and hence a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra as explained in 6.4).
-We will show this by constructing an explicit $\operatorname{\mathbb{F}}_{q}$-linear functional on $A.$ We learned about this construction in [19].
-Every element $g$ in $\operatorname{\mathbb{F}}_{q}[x]/f$ can be written uniquely in the form
-$g=a_{0}(g)+a_{1}(g)+\cdots+a_{n-1}(g)x^{n-1}.$
-We define $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ by
+Let $A$ be a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra together with a non-degenerate functional $\lambda\colon A\to\operatorname{\mathbb{F}}_{p}.$ For each $a\in A$, let $\lambda_{a}\colon A\to\operatorname{\mathbb{C}}^{\times}$ defined by $\lambda_{a}(b)=\zeta_{p}^{\lambda(ab)}.$ Then $\lambda_{a}\in\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times}).$ Furthermore, the map $a\mapsto\lambda_{a}$ gives an isomorphism between $A$ and $\operatorname{{\rm Hom}}((A,+),\operatorname{\mathbb{C}}^{\times})$.
+We will now focus on the case $A=\operatorname{\mathbb{F}}_{q}[x]/(f)$.
+We will show that it is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra (and hence a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra as explained from 6.4).
+We will show this by constructing an explicit $\operatorname{\mathbb{F}}_{q}$-linear functional on $A.$ We learned about this construction from [20].
+Assume $\deg f=n$.
+Every element $g$ in $\operatorname{\mathbb{F}}_{q}[x]/(f)$ can be written uniquely in the form
+$g=a_{0}(g)+a_{1}(g)x+\cdots+a_{n-1}(g)x^{n-1}.$
+We define $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ by
 $\psi(g)=a_{n-1}(g).$
 Proposition 6.7.
-Suppose $g\in\operatorname{\mathbb{F}}_{q}[x]/f$ such that $\psi(hg)=0$ for all $h\in\operatorname{\mathbb{F}}_{q}[x]/f.$ Then $g=0$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$
+Suppose $g\in\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $\psi(hg)=0$ for all $h\in\operatorname{\mathbb{F}}_{q}[x]/(f).$ Then $g=0$ in $\operatorname{\mathbb{F}}_{q}[x]/(f).$
 Proof.
-Let us write $g=a_{0}(g)+a_{1}(g)+\cdots+a_{n-1}(g)x^{n-1}$.
-We will prove by induction that $a_{n-k}(g)=0$ for $1\leq k\leq n.$ In fact, since $\psi(g)=0$, we know that $a_{n-1}(g)=0.$ Consequently, the statement is true for $k=1.$ Let us assume that it has been for all $1\leq k\leq m<n$.
+Let us write $g=a_{0}(g)+a_{1}(g)x+\cdots+a_{n-1}(g)x^{n-1}$.
+We will prove by induction that $a_{n-k}(g)=0$ for $1\leq k\leq n.$ In fact, since $\psi(g)=0$, we know that $a_{n-1}(g)=0.$ Consequently, the statement is true for $k=1.$ Let us assume that it has been shown for all $1\leq k\leq m<n$.
 We claim that it is also true for $m+1$, namely $a_{n-m-1}(g)=0$ as well.
 In fact, we have
-$\begin{split}x^{m-1}g&=x^{m-1}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}+a_{n-m}(g)x^{n}+\cdots+a_{n-1}(g)x^{m+n-2}\\ &=x^{m-1}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}.\end{split}$
-Consequently
-$a_{n-m-1}(g)=\psi(x^{m-1}g)=0.$
+$\begin{split}x^{m}g&=x^{m}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}+a_{n-m}(g)x^{n}+\cdots+a_{n-1}(g)x^{m+n-1}\\ &=x^{m}a_{0}(g)+\cdots+a_{n-m-1}(g)x^{n-1}.\end{split}$
+Here, the last equation follows from the fact that $a_{n-m}=a_{n-m+1}=\ldots=a_{n-1}=0$ by the induction hypothesis. Consequently
+$a_{n-m-1}(g)=\psi(x^{m}g)=0.$
 By the induction principle, we conclude that $a_{n-k}=0$ for all $1\leq k\leq n.$
 Corollary 6.8.
-$\psi$ is a non-degenerate $\operatorname{\mathbb{F}}_{q}$-linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f.$ Consequently, $\operatorname{\mathbb{F}}_{q}[x]/f$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra.
+$\psi$ is a non-degenerate $\operatorname{\mathbb{F}}_{q}$-linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f).$ Consequently, $\operatorname{\mathbb{F}}_{q}[x]/(f)$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra.
+By 6.4, under the composition $\operatorname{\mathbb{F}}_{q}[x]/(f)\xrightarrow{\psi}\operatorname{\mathbb{F}}_{q}\xrightarrow{\operatorname{{\rm Tr}}}\operatorname{\mathbb{F}}_{p}$ where Tr is the trace map, $\operatorname{\mathbb{F}}_{q}[x]/(f)$ becomes a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra.
+By 6.6, we have the following proposition.
+Proposition 6.9.
+There exists a bijection
+$\operatorname{\mathbb{F}}_{q}[x]/(f)\longleftrightarrow\operatorname{{\rm Hom}}((\operatorname{\mathbb{F}}_{q}[x]/(f),+),\operatorname{\mathbb{C}}^{\times}),\quad a\longleftrightarrow\{\psi_{a}\},$
+where
+$\psi_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{C}}^{\times}$
+is given by
+$\psi_{a}(b)=\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ab))},\forall b\in\operatorname{\mathbb{F}}_{q}[x]/(f).$
 […]
 Proposition 6.10.
-There exists a bijection
-$\operatorname{\mathbb{F}}_{q}[x]/f\longleftrightarrow\operatorname{{\rm Hom}}((\operatorname{\mathbb{F}}_{q}[x]/f,+),\operatorname{\mathbb{C}}^{\times}),\quad a\longleftrightarrow\{\psi_{a}\},$
-where
-$\psi_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{C}}^{\times}$
-is given by
-$\psi_{a}(b)=\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ab))},\forall b\in\operatorname{\mathbb{F}}_{q}[x]/f.$
-We remark that over $\operatorname{\mathbb{Z}}$, if $\zeta_{n}$ is a primitive $n$th-root of unity, then for each divisor $m\mid n$, $\zeta_{n}^{\frac{n}{m}}$ is a primitive $m$-root of unity.
-An analogous statement holds for $\operatorname{\mathbb{F}}_{q}[x]$ as well.
-Proposition 6.10.
-Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional.
+Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional.
 Let $g$ be a divisor of $f$ and $\psi_{g}\colon\operatorname{\mathbb{F}}_{q}[x]/g\to\operatorname{\mathbb{F}}_{q}$ be the function defined by
@@
 We only need to show that it is non-degenerate.
-In fact, suppose to the contrary that the kernel of $\psi_{g}$ contains a non-zero ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ Since $\operatorname{\mathbb{F}}_{q}[x]$ is a PID, $I$ must be of the form $I=\langle h\rangle$ for some $h|g.$ We then see that $\langle h\frac{f}{g}\rangle$ belongs to the kernel of $\psi.$ Because $\psi$ is non-degenerate, this implies that $h\frac{f}{g}=0$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$ In other words, $g|h$ or equivalently $h=0$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ This shows that $I=0$, which is a contradiction.
+In fact, suppose to the contrary that the kernel of $\psi_{g}$ contains a non-zero ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ Since $\operatorname{\mathbb{F}}_{q}[x]$ is a PID, $I$ must be of the form $I=\langle h\rangle$ for some $h\mid g.$ We then see that $\langle h\frac{f}{g}\rangle$ belongs to the kernel of $\psi.$ Because $\psi$ is non-degenerate, this implies that $h\frac{f}{g}=0$ in $\operatorname{\mathbb{F}}_{q}[x]/(f).$ In other words, $g\mid h$ or equivalently $h=0$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ This shows that $I=0$, which is a contradiction.
 ∎
@@
 Definition 6.11.
-(Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$) Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be monic polynomials.
-Let $\psi:\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f$.
+(Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$) Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be polynomials such that $f$ is monic.
+Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f)$.
 The Ramanujan sum $c(g,f)$ is defined as
-$c(g,f)=c_{\psi}(g,f)=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}.$
+$c(g,f)=c_{\psi}(g,f)=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}.$
 Remark 6.12.
-We remark also that at first glance, $c(g,f)$ depends on the choice of $\psi.$ However, as we explain in what follows, it does not as long as we make sure that $\psi$ is non-degenerate.
-This is similar to the case over $\operatorname{\mathbb{Z}}$: Ramanujan sums do not depend on the choice of a primitive $n$th-root of unity.
-In fact, we will show that there is an explicit formula for $c(g,f)$ similar to the case of Ramanujan sum over $\operatorname{\mathbb{Z}}.$
+[…]
 Remark 6.13.
-Ramanujan sums are a special case of Gauss sums as defined and studied in [21, Definition 1].
-In fact, they are Gauss sums for the principal Dirichlet characters on $\operatorname{\mathbb{F}}_{q}[x]/f.$ It would be rather interesting if we could define Fekete polynomials for these principal Dirichlet characters (the case over $\operatorname{\mathbb{Z}}$ was studied in [7]).
+[…]
 We first recall the following standard lemma in group theory.
@@
 Because $\chi(h)\neq 1$, we must have $\sum_{g\in G}\chi(g)=0.$
+The Möbius function $\mu$ on $\operatorname{\mathbb{F}}_{q}[x]$ is defined in the same way as in the integer case, which we recall as follows: for a non-zero polynomial $f\in\operatorname{\mathbb{F}}_{q}[x]$,
+$\mu(f)=\begin{cases}(-1)^{r}&\text{if $f=aP_{1}\cdots P_{r}$, }\\ 0&\text{ if $f$ is not square-free.}\end{cases}$
+Here $a$ is a non-zero constant and $P_{1},\ldots,P_{r}$ are distinct monic irreducible polynomials dividing $f$ ($r$ could be 0).
 Proposition 6.15.
-Let $\psi$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f$. For each $f\in\operatorname{\mathbb{F}}_{q}[x]$
+Let $\psi$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f)$. For each $f\in\operatorname{\mathbb{F}}_{q}[x]$
 $c(1,f)=c_{\psi}(1,f)=\mu(f),$
@@
 Proof.
-If $f=\prod_{i=1}^{d}f_{i}^{\alpha_{i}}$ where $\alpha_{i}\in\operatorname{\mathbb{N}}$ and $f_{i}$ are irreducible, then by the Chinese remainder theorem and [21, Satz 1] we have
+If $f=\prod_{i=1}^{d}f_{i}^{\alpha_{i}}$ where $\alpha_{i}\in\operatorname{\mathbb{N}}$ and $f_{i}$ are irreducible, then by the Chinese remainder theorem and [22, Satz 1] we have
 $c_{\psi}(1,f)=\prod_{i=1}^{d}c_{\psi_{i}}(1,f_{i}^{\alpha_{i}}),$
-where $\psi_{i}:\operatorname{\mathbb{F}}_{q}[x]/f_{i}^{\alpha_{i}}\to\operatorname{\mathbb{F}}_{q}$ is a non-degenerate linear functional.
+where $\psi_{i}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)_{i}^{\alpha_{i}}\to\operatorname{\mathbb{F}}_{q}$ is a non-degenerate linear functional induced by $\psi.$
 Therefore, it is enough to consider the case $f=f_{1}^{a_{1}}$ where $f_{1}$ is irreducible. We have
@@
 Similarly, if $a_{1}\geq 2$ then by 6.14 and 6.10
-$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=0.$
-On the other hand if $a_{1}=1$ then $\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=-1$. We conclude that
+$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/(f_{1}^{a_{1}-1})}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=0.$
+On the other hand if $a_{1}=1$ then $\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/(f_{1}^{a_{1}-1})}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=1$. We conclude that
 $c(1,f)=c_{\psi}(1,f)=c_{\psi}(1,f_{1}^{a_{1}})=\mu(f_{1}^{a_{1}})=\mu(f).\qed$
@@
 Theorem 6.16.
-Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be monic polynomials. Then (compare with [18, Equation 9])
+Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be polynomials such that $f$ is monic. Then (compare with [19, Equation 9])
 $c(g,f)=\mu(t)\dfrac{\varphi(f)}{\varphi(t)},\quad\text{where}\quad t=\dfrac{f}{\gcd(f,g)}.$
+Here $\varphi$ is the Euler totient function on $\operatorname{\mathbb{F}}_{q}[x]$; i.e, $\varphi(h)=|(\operatorname{\mathbb{F}}_{q}[x]/(h))^{\times}|.$ In particular, $c(g,f)$ is an integer.
 Proof.
-Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional and
+Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional and
 $c(g,f)=c_{\psi}(g,f).$
-Let $h=\gcd(g,f)$ and $t=\dfrac{f}{h}$.
-The canonical projection map $(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}\to(\operatorname{\mathbb{F}}_{q}[x]/t)^{\times}$ has kernel of size $\frac{\varphi(f)}{\varphi(t)}.$ Furthermore, for each $a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$, if $a_{1},a_{2}$ are two preimages of $a$ in $(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}$ then $a_{2}=a_{1}+tu$, for some $u\in\operatorname{\mathbb{F}}_{q}[x]/f$.
-In this case, one has $\psi(ga_{2})=\psi(ag_{1}+gtu)=\psi(ga_{1})+\psi(\frac{gu}{h}f)=\psi(ga_{1})$.
+Let $h=\gcd(g,f)$ and $t=\dfrac{f}{h}$. The canonical projection map
+$(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}\to(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$},
+has kernel of size $\frac{\varphi(f)}{\varphi(t)}.$ Furthermore, for each $a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$, if $a_{1},a_{2}$ are two preimages of $a$ in $(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}$ then $a_{2}=a_{1}+tu$, for some $u\in\operatorname{\mathbb{F}}_{q}[x]/(f)$.
+In this case, one has
+$\psi(ga_{2})=\psi(ag_{1}+gtu)=\psi(ga_{1})+\psi(\frac{gu}{h}f)=\psi(ga_{1}).$
 Consequently, we have
-$c_{\psi}(g,f)$ $=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/t)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}$
+$c_{\psi}(g,f)$ $=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}$
 $=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(\frac{f}{t}\frac{g}{h}a))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{t}(\frac{g}{h}a))}.$
-Here $\psi_{t}:\operatorname{\mathbb{F}}_{q}[x]/t\to\operatorname{\mathbb{F}}_{q}$ is the linear functional defined by $\psi_{t}(b)=\psi(\frac{f}{t}b)=\psi(hb)$.
+Here $\psi_{t}\colon\operatorname{\mathbb{F}}_{q}[x]/(t)\to\operatorname{\mathbb{F}}_{q}$ is the linear functional defined by $\psi_{t}(b)=\psi(\frac{f}{t}b)=\psi(hb)$.
 Note that $\psi_{t}$ is non-degenerate.
@@
 The adjacency matrix of the Cayley graph $\Gamma(G,S)$ is a $G$-circulant matrix.
-By the $G$-circulant theorem, the spectrum of $\Gamma(G,S)$ is the collection of the sums $\sum_{s\in S}\chi(s)$, where $\chi$ runs over the set of all characters of $G$ (see [14, Section 1.2]).
-When $G=\operatorname{\mathbb{F}}_{q}[x]/f$, we know by 6.9 that the character of $G$ is parameterized by $G$ itself.
+By the $G$-circulant theorem, the spectrum of $\Gamma(G,S)$ is the collection of the sums $\sum_{s\in S}\chi(s)$, where $\chi$ runs over the set of all characters of $G$ (see [15, Section 1.2, Equation (1.1’)]).
+When $G=\operatorname{\mathbb{F}}_{q}[x]/(f)$, we know by 6.9 that the character of $G$ is parameterized by $G$ itself.
 Consequently, we have the following proposition.
 Proposition 6.17.
-Let $S$ be a subset of $\operatorname{\mathbb{F}}_{q}[x].$ Then the spectrum of the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ is given by the set
-$\left\{\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/f}.$
+Let $S$ be a subset of $\operatorname{\mathbb{F}}_{q}[x].$ Then the eigenvalues of the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ are given by the set
+$\left\{\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/(f)}.$
 Let us now focus on the case of gcd-graphs; i.e., $S=S_{D}$. We have the following lemma.
@@
 Let $h\mid f$ be a monic divisor of $f.$ Then for each $g\in\operatorname{\mathbb{F}}_{q}[x]$
-$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f,\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=c\left(g,\frac{f}{h}\right).$
+$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/(f),\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=c\left(g,\frac{f}{h}\right).$
 Proof.
 For $a\in\operatorname{\mathbb{F}}_{q}[x]$, $\gcd(a,f)=h$ if and only if $a=hb$ where $\gcd(b,\frac{f}{h})=1.$ Therefore, the above sum can be rewritten as
-$\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(a,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(bgh))}=\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(b,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}.$
-Here $\psi_{h}:\operatorname{\mathbb{F}}_{q}[x]/(f/h)\to\operatorname{\mathbb{F}}_{q}$ is the functional given by $\psi_{h}(x)=\psi(hx).$ By 6.10, $\psi_{h}$ is a non-generate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f/g).$ Therefore, we conclude that
-$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f,\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(b,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}=c\left(g,\frac{f}{h}\right).$
+$\sum_{\begin{subarray}{c}b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h)\\ \gcd(a,f/h)=1\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(bgh))}=\sum_{\begin{subarray}{c}b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h)\\ \gcd(b,f/h)=1\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}.$
+Here $\psi_{h}\colon\operatorname{\mathbb{F}}_{q}[x]/(f/h)\to\operatorname{\mathbb{F}}_{q}$ is the functional given by $\psi_{h}(x)=\psi(hx).$ By 6.10, $\psi_{h}$ is a non-generate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f/g).$ Therefore, we conclude that
+$\sum_{\begin{subarray}{c}a\in\operatorname{\mathbb{F}}_{q}[x]/(f)\\ \gcd(a,f)=h\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=\sum_{\begin{subarray}{c}b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h)\\ \gcd(b,f/h)=1\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}=c\left(g,\frac{f}{h}\right).$
 By 6.17 and 6.18, we have the following theorem.
 Theorem 6.19.
-Let $f$ be a monic polynomial and $D=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}\mid f.$ Then, the spectrum of $S_{f}(D)$ is given by the set
-$\left\{\sum_{i=1}^{k}c\left(g,\frac{f}{f_{i}}\right)\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/f}.$
-Corollary 6.20.
-All eigenvalues of the gcd-graph $G_{f}(D)$ are integers.
-As we discussed in the introduction, it is known that a $\operatorname{\mathbb{Z}}/n$-circulant graph has an integral spectrum if and only if it is a gcd-graph (see [29, Theorem 7.1] ).
-One may ask whether the same statement is true for graphs associated with $\operatorname{\mathbb{F}}_{q}[x]/f.$ The answer is no in general.
+Let $f$ be a monic polynomial and $D=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}\mid f.$ Then, the eigenvalues of $G_{f}(D)$ are given by the set
+$\left\{\sum_{i=1}^{k}c\left(g,\frac{f}{f_{i}}\right)\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/(f)}.$
+In particular, all eigenvalues of the gcd-graph $G_{f}(D)$ are integers.
+Example 6.20.
+The characteristic polynomial of the graph described in Fig. 1 has the following factorization
+$(x-4)(x-1)^{4}(x+2)^{4}.$
+Similarly, the characteristic polynomial of the graph described in Fig. 2 has the following factorization
+$(x-3)(x+3)(x-1)^{3}(x+1)^{3}.$
+In particular, the eigenvalues of these graphs are integers as shown in 6.19.
+We remark that these spectral calculations have also been verified using Ramanujan sums, as explained in 6.19.
+These verifications are described in [24], which contains the code to calculate various arithmetic functions over $\operatorname{\mathbb{F}}_{q}[x]$.
+In particular, it has a function to calculate the general Ramanujan sum $c(g,f)$ for all $g,f\in\operatorname{\mathbb{F}}_{q}[x].$
+For example, let us consider the graph $G_{f}(D)$ described in Fig.
+1, where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$} and $D=\{x,x+1\}$.
+We shall identify $\operatorname{\mathbb{F}}_{3}[x]/(f)$ with
+$\{0,1,2,x,x+1,x+2,2x,2x+1,2x+2\}.$
+For $g=0$, we have $c(0,x)=\mu(1)\dfrac{\varphi(x)}{\varphi(1)}=2$, and similarly $c(0,x+1)=2$.
+Hence the eigenvalue of $G_{f}(D)$ associated with $g=0$ is $c(0,x)+c(0,x+1)=4$.
+(The eigenvalue $4$ is the degree of $G_{f}(D)$, and it appears with multiplicity $1$.)
+For $g$ coprime to $x$, we have $c(g,x)=\mu(x)\dfrac{\varphi(x)}{\varphi(x)}=-1$. Similarly, for $g$ coprime to $x+1$, we have c(g,x+1)=1c(g,x+1)=-1. We also have
+$c(x,x)=c(2x,x)=\mu(1)\dfrac{\varphi(x)}{\varphi(1)}=2,$
+and
+$c(x+1,x+1)=c(2x+2,x+1)=2.$
+From the above discussion, we see that if $g\in\{1,2,x+2,2x+1\}$, then the eigenvalue of $G_{f}(D)$ associated with $g$ is
+c(g,x)+c(g,x+1)=(1)+(1)=2.c(g,x)+c(g,x+1)=(-1)+(-1)=-2.
+(The eigenvalue 2-2 appears with multiplicity 4.) Finally, if $g\in\{x,x+1,2x,2x+2\}$, then the eigenvalue of $G_{f}(D)$ associated with $g$ is
+c(g,x)+c(g,x+1)=2+(1)=1.c(g,x)+c(g,x+1)=2+(-1)=1.
+(The eigenvalue $1$ appears with multiplicity 4.) Our calculations show that the spectrum of $G_{f}(D)$, as described in 6.19, coincides with the spectrum obtained from the adjacency matrix.
+As we discussed in the introduction, it is known that a $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$-circulant graph has an integral spectrum if and only if it is a gcd-graph (see [32, Theorem 7.1] ).
+One may ask whether the same statement is true for graphs associated with $\operatorname{\mathbb{F}}_{q}[x]/(f).$ The answer is no in general.
 In fact, we have the following general observation.
 Proposition 6.21.
-Let $S$ be a symmetric subset of $\operatorname{\mathbb{F}}_{q}[x]/f$ such that $0\not\in S.$ Suppose further that $\operatorname{\mathbb{F}}_{p}^{\times}S=S.$ Then, the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ has an integral spectrum.
+Let $S$ be a symmetric subset of $\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $0\not\in S.$ Suppose further that $\operatorname{\mathbb{F}}_{p}^{\times}S=S.$ Then, the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ has an integral spectrum.
 Proof.
-We define the following equivalence relation on $\operatorname{\mathbb{F}}_{q}[x]/f$.
-For $u,v\in\operatorname{\mathbb{F}}_{q}[x]/f$, we say that $u\sim v$ if $u=kv$ where $k\in\operatorname{\mathbb{F}}_{p}^{\times}.$ By 6.17, it is enough to show that for each $g\in\operatorname{\mathbb{F}}_{q}[x]/f$, $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$.
+We define the following equivalence relation on $\operatorname{\mathbb{F}}_{q}[x]/(f)$.
+For $u,v\in\operatorname{\mathbb{F}}_{q}[x]/(f)$, we say that $u\sim v$ if $u=kv$ where $k\in\operatorname{\mathbb{F}}_{p}^{\times}.$ By 6.17, it is enough to show that for each $g\in\operatorname{\mathbb{F}}_{q}[x]/(f)$, $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$.
 Because $\operatorname{\mathbb{F}}_{p}^{\times}S=S$, this sum can be written as
@@
 $\sum_{k\in\operatorname{\mathbb{F}}_{p}^{\times}}(\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))})^{k}=\begin{cases}-1&\text{if }\operatorname{{\rm Tr}}(\psi(gs))\neq 0\\ p-1&\text{if }\operatorname{{\rm Tr}}(\psi(gs))=0.\end{cases}$
-This shows that $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$ for each $g\in\operatorname{\mathbb{F}}_{q}[x]/f.$
+This shows that $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$ for each $g\in\operatorname{\mathbb{F}}_{q}[x]/(f).$
 Remark 6.22.
 Let $\operatorname{\mathbb{F}}_{q}$ be a finite field such that $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{p}.$ Let $f=x$.
-In this case $\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{q}.$ Let $V$ be proper $\operatorname{\mathbb{F}}_{p}$-subspace of $\operatorname{\mathbb{F}}_{q}$ and $S=V\setminus\{0\}.$ Then by 6.21, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ has an integral spectrum even though it is not a gcd-graph.
-We wonder whether the converse of 6.21 holds (perhaps, under some mild assumptions).
+In this case $\operatorname{\mathbb{F}}_{q}[x]/(f)\cong\operatorname{\mathbb{F}}_{q}.$ Let $V$ be proper $\operatorname{\mathbb{F}}_{p}$-subspace of $\operatorname{\mathbb{F}}_{q}$ and $S=V\setminus\{0\}.$ Then by 6.21, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ has an integral spectrum even though it is not a gcd-graph.
+In a recent work, we show that the converse of this statement is also true; namely, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ is integral if and only if $S\cup\{0\}$ is a vector space over $\operatorname{\mathbb{F}}_{p}$ (see [28, Theorem 1.1].)
 7. Perfect gcd-graphs
@@
 In [1, Theorem 9.5], the authors classify all perfect unitary Cayley graphs associated with finite commutative rings.
-Specifically, for a ring $R=R_{1}\times R_{2}\times\cdots\times R_{t}$ where $R_{i}$ are finite local rings such that $|R_{1}|\leq|R_{2}|\leq\ldots\leq|R_{t}|$, the unitary Cayley graph on $R$ is perfect if and only if one of the following conditions hold
+Specifically, for a ring $R=R_{1}\times R_{2}\times\cdots\times R_{t}$ where $R_{i}$ are finite local rings such that $|R_{1}|\leq|R_{2}|\leq\cdots\leq|R_{t}|$, the unitary Cayley graph on $R$ is perfect if and only if one of the following conditions hold:
   1. (1)
-    The residue field of $R_{1}$ is $2$. In this case, $G_{R}$ is bipartite and hence perfect.
+    The residue field of $R_{1}$ is $2$. In this case, $G_{R}$ is bipartite and hence perfect;
   2. (2)
     $R$ is either a local ring or a product of two local rings; i.e, $t\leq 2.$
-A direct consequence of [1, Theorem 9.5] for the unitary Cayley graph over $\operatorname{\mathbb{Z}}$ is the following.
+[…]
 Corollary 7.1.
-The unitary Cayley graph on $\operatorname{\mathbb{Z}}/N$ is perfect if and only if one of the following conditions holds.
+The unitary Cayley graph on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ is perfect if and only if one of the following conditions holds.
   1. (1)
@@
 Corollary 7.2.
-The unitary Cayley graph $G_{f}(\{1\})$ on $\operatorname{\mathbb{F}}_{q}[x]/f$ is perfect if and only if one of the following conditions holds.
+The unitary Cayley graph $G_{f}(\{1\})$ on $\operatorname{\mathbb{F}}_{q}[x]/(f)$ is perfect if and only if one of the following conditions holds.
   1. (1)
@@
 We will use 7.2 together with patterns of $G_{f}(D)$ discovered through our experiments in Sagemath and the Python package Networkx in order to find some sufficient conditions for $G_{f}(D)$ to be non-perfect.
-In particular, we will exploit the fact that certain induced subgraphs of $G_{f}(D)$ are naturally isomorphic to the unitary Cayley graphs on some quotient rings of $R=\operatorname{\mathbb{F}}_{q}[x]/f.$ More precisely, by 5.4, we have the following observation.
+In particular, we will exploit the fact that certain induced subgraphs of $G_{f}(D)$ are naturally isomorphic to the unitary Cayley graphs on some quotient rings of $R=\operatorname{\mathbb{F}}_{q}[x]/(f).$ More precisely, by 5.4, we have the following observation.
 Proposition 7.3.
@@
 In this case
-$\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{q}[x]/f_{1}\times\operatorname{\mathbb{F}}_{q}[x]/f_{2}\times\operatorname{\mathbb{F}}_{q}[x]/f_{3}.$
-Under this isomorphism, we can identify $V(G_{f}(D))$ as the set of all triples $(a_{1},a_{2},a_{3})$ where $a_{i}\in\operatorname{\mathbb{F}}_{q}[x]/f_{i}$ for $1\leq i\leq 3$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if there exists an index $i\in\{1,2,3\}$ such that $a_{i}=b_{i}$ and $(a_{j}-b_{i})\in(\operatorname{\mathbb{F}}_{q}[x]/f_{j})^{\times}$ if $j\neq i$.
+$\operatorname{\mathbb{F}}_{q}[x]/(f)\cong\operatorname{\mathbb{F}}_{q}[x]/(f_{1})\times\operatorname{\mathbb{F}}_{q}[x]/(f_{2})\times\operatorname{\mathbb{F}}_{q}[x]/(f_{3}).$
+Under this isomorphism, we can identify $V(G_{f}(D))$ as the set of all triples $(a_{1},a_{2},a_{3})$ where $a_{i}\in\operatorname{\mathbb{F}}_{q}[x]/(f_{i})$ for $1\leq i\leq 3.$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if there exists an index $i\in\{1,2,3\}$ such that $a_{i}=b_{i}$ and $(a_{j}-b_{i})\in(\operatorname{\mathbb{F}}_{q}[x]/(f_{j}))^{\times}$} if $j\neq i$.
 Using Sagemath, we can find the following induced $5$-cycle in $G_{f}(D)$
@@
 If $\omega(f)>\omega(f_{1})+\omega(f_{2})$ then $\omega(f/f_{2})>\omega(f_{1})\geq 2.$ By 7.3, we know $G_{f}(D)$ is not perfect. Therefore, we must have $\omega(f)=\omega(f_{1})+\omega(f_{2}).$
-Let $g=f_{1}f_{2}.$ By 5.6, the ideal generated by $g$ is a homogeneous set in $G_{f}(D)$ and furthermore $G_{f}(D)$ is isomorphic to the wreath product $G_{g}(\{f_{1},f_{2}\})*G_{f/g}(\emptyset).$ Since $G_{f}(D)$ is perfect, $G_{g}(\{f_{1},f_{2}\})$ is perfect as well.
+Let $g=f_{1}f_{2}.$ By 5.6, the ideal generated by $g$ is a homogeneous set in $G_{f}(D)$ and furthermore $G_{f}(D)$ is isomorphic to the wreath product $G_{g}(\{f_{1},f_{2}\})*G_{f/g}(\emptyset)$}.
+Since $G_{f}(D)$ is perfect, $G_{g}(\{f_{1},f_{2}\})$ is perfect as well.
 Because $\omega(f_{1})\geq 2$, we can find $h_{1},h_{2}\in\operatorname{\mathbb{F}}_{q}[x]$ such that $f_{1}=h_{1}h_{2}$ and $\gcd(h_{1},h_{2})=1.$ By the Chinese remainder theorem
-$\operatorname{\mathbb{F}}_{q}[x]/g\cong\operatorname{\mathbb{F}}_{q}[x]/h_{1}\times\operatorname{\mathbb{F}}_{q}[x]/h_{2}\times\operatorname{\mathbb{F}}_{q}[x]/f_{2}.$
-Under this isomorphism, we can identify $\operatorname{\mathbb{F}}_{q}[x]/g$ with the set of tuples $(a_{1},a_{2},a_{3})$ such that $a_{1}\in\operatorname{\mathbb{F}}_{q}[x]/h_{1},a_{2}\in\operatorname{\mathbb{F}}_{q}[x]/h_{2},a_{3}\in\operatorname{\mathbb{F}}_{q}[x]/f_{2}.$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if one of the following conditions happens
+$\operatorname{\mathbb{F}}_{q}[x]/(g)\cong\operatorname{\mathbb{F}}_{q}[x]/(h_{1})\times\operatorname{\mathbb{F}}_{q}[x]/(h_{2})\times\operatorname{\mathbb{F}}_{q}[x]/(f_{2}).$
+Under this isomorphism, we can identify $\operatorname{\mathbb{F}}_{q}[x]/(g)$ with the set of tuples $(a_{1},a_{2},a_{3})$ such that $a_{1}\in\operatorname{\mathbb{F}}_{q}[x]/(h_{1}),a_{2}\in\operatorname{\mathbb{F}}_{q}[x]/(h_{2}),a_{3}\in\operatorname{\mathbb{F}}_{q}[x]/(f_{2}).$ Furthermore, two vertices $(a_{1},a_{2},a_{3})$ and $(b_{1},b_{2},b_{3})$ are adjacent if and only if one of the following conditions happens
   1. (1)
-    $a_{1}=b_{1},a_{2}=b_{2}$ and $a_{3}-b_{3}\in(\operatorname{\mathbb{F}}_{q}[x]/f_{2})^{\times}$,
+    $a_{1}=b_{1},a_{2}=b_{2}$ and $a_{3}-b_{3}\in(\operatorname{\mathbb{F}}_{q}[x]/(f_{2}))^{\times}$,
   2. (2)
-    $a_{3}=b_{3}$ and $a_{i}-b_{i}\in(\operatorname{\mathbb{F}}_{q}[x]/h_{i})^{\times}$ for $i\in\{1,2\}.$
+    $a_{3}=b_{3}$ and $a_{i}-b_{i}\in(\operatorname{\mathbb{F}}_{q}[x]/(h_{i}))^{\times}$ for $i\in\{1,2\}.$
 Using Sagemath, we can find the following $7$-cycle in $G_{g}(\{f_{1},f_{2}\})$
@@
 Curious readers might wonder why we choose a $7$-cycle in the proof of 7.5 instead of choosing a $5$-cycle as in the proof of 7.4.
-The reason is that when $\omega(f_{1})=2,\omega(f_{2})=1$ and $f=f_{1}f_{2}$, our code cannot find $5$-cycle in $G_{f}(D).$ It seems interesting to investigate whether this is always the case.
+The reason is that when $\omega(f_{1})=2,\omega(f_{2})=1$ and $f=f_{1}f_{2}$, our code cannot find a $5$-cycle in $G_{f}(D).$ It seems interesting to investigate whether this is always the case.
 Remark 7.7.
 The only remaining case that we miss is when $\omega(f_{1})=\omega(f_{2})=1.$ We distinguish the following cases.
-  1. (1)
-    If $f=f_{1}f_{2}$ where both $f_{1},f_{2}$ are irreducible then $G_{f}(D)$ is isomorphic to the complement of $G_{f}(\{1\})$ which is perfect by 7.2. Therefore, $G_{f}(D)$ is perfect as well.
-  2. (2)
-    On the other hand, if either $f_{1}$ or $f_{2}$ is reducible, our code can always find a $7$-cycle in $G_{f}(D).$ Unfortunately, we cannot find a universal pattern in this case.
+(1) If $f=f_{1}f_{2}$ where both $f_{1},f_{2}$ are irreducible then $G_{f}(D)$ is isomorphic to the complement of $G_{f}(\{1\})$ which is perfect by 7.2. Therefore, $G_{f}(D)$ is perfect as well.
+(2) On the other hand, if either $f_{1}$ or $f_{2}$ is reducible, our code can always find a $7$-cycle in $G_{f}(D).$ Unfortunately, we cannot find a universal pattern in this case.
 It would be quite interesting to solve this puzzle completely.
 8. Induced subgraphs of gcd-graphs
-A theorem of Erdős and Evans (see [11]) says that every graph $G$ is an induced subgraph of the unitary Cayley graph on $\operatorname{\mathbb{Z}}/n$ for some squarefree $n.$ Using this result, it is shown in [16] that for a given finite graph $G$ and a finite field $F$, $G$ is an induced subgraph of the unitary graph of a matrix algebra $M_{d}(F)$ for some value of $d$ (they also provide some precise upper-bound on $d$ when $G$ is the complete graph $K_{m}$).
+A theorem of Erdős and Evans (see [12]) says that every graph $G$ is an induced subgraph of the unitary Cayley graph on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$ for some squarefree $n.$ Using this result, it is shown in [17] that for a given finite graph $G$ and a finite field $F$, $G$ is an induced subgraph of the unitary graph of a matrix algebra $M_{d}(F)$ for some value of $d$ (they also provide some precise upper-bound on $d$ when $G$ is the complete graph $K_{m}$).
 In light of these results, it seems interesting to ask whether a graph $G$ can be realized as an induced subgraph of $G_{f}(D)$ for some choice of $f,D$ and $\operatorname{\mathbb{F}}_{q}.$ Under some rather mild conditions, the answer is yes as we will show below.
+For the related definitions we refer the reader to Section 2.
 First, we introduce the following observation.
 Lemma 8.1.
-If $G$ is an induced subgraph of $H$, then $G$ is also an induced subgraph of the product $H\times K_{|G|}$ where $K_{|G|}$ is the complete graph on $|G|$-nodes.
+If $G$ is an induced subgraph of $H$, then $G$ is also an induced subgraph of the tensor product $H\times K_{|G|}$ where $K_{|G|}$ is the complete graph on $|G|$-nodes.
 Proof.
@@
 Let $G$ be a fixed graph and $q$ a fixed prime power.
-There exists a positive integer $r$ such that for each $d\geq r$, we can find $F_{d}\in\operatorname{\mathbb{F}}_{q}[x]$ satisfying the following conditions
+There exists a positive integer $r$ such that for each $d\geq r$, there exists $F_{d}\in\operatorname{\mathbb{F}}_{q}[x]$ satisfying the following conditions
   1. (1)
@@
 Proof.
-By Erdős and Evans’s theorem [11], there exists a squarefree integer $n\in\operatorname{\mathbb{N}}$ such that $G$ is an induced subgraph of the unitary graph $G_{n}$ on $\operatorname{\mathbb{Z}}/n.$ Let $r=\omega(n)$; i.e, $n=p_{1}p_{2}\ldots p_{r}$ be the prime factorization of $n$.
+By Erdős and Evans’s theorem [12], there exists a squarefree integer $n\in\operatorname{\mathbb{N}}$ such that $G$ is an induced subgraph of the unitary graph $G_{n}$ on $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}.$ Let $r=\omega(n)$; i.e, $n=p_{1}p_{2}\ldots p_{r}$ be the prime factorization of $n$.
 Then
 $G_{n}\cong\prod_{i=1}^{r}G_{p_{i}}\cong\prod_{i=1}^{r}K_{p_{i}}.$
-For each $i\in\operatorname{\mathbb{N}}$ and $m_{i}\in\operatorname{\mathbb{N}}$, there exists a polynomial $h_{i}$ of degree $m_{i}$ such that $h_{i}$ is irreducible over $\operatorname{\mathbb{F}}_{q}[x].$ We then see that $G_{h_{i}}(\{1\})$ is isomorphic to the complete graph $K_{q^{m_{i}}}.$ Let us choose $m_{i}$ so that $q^{m_{i}}>n$ for all $i$.
+For each $i\in\operatorname{\mathbb{N}}$ and $m_{i}\in\operatorname{\mathbb{N}}$, by Galois theory for finite fields, there exists a polynomial $h_{i}$ of degree $m_{i}$ such that $h_{i}$ is irreducible over $\operatorname{\mathbb{F}}_{q}[x].$ We then see that $G_{h_{i}}(\{1\})$ is isomorphic to the complete graph $K_{q^{m_{i}}}.$ Let us choose $m_{i}$ so that $q^{m_{i}}>n$ for all $i$.
 For each $d\geq r$, let $F_{d}=h_{1}h_{2}\ldots h_{d}.$ We then have
 $G_{F_{d}}(\{1\})\cong\prod_{i=1}^{d}G_{h_{i}}(\{1\})\cong\prod_{i=1}^{d}K_{q^{m_{i}}}.$
-Since $q^{m_{i}}>n$, we know that $G_{n}$ is an induced subgraph of $G_{F_{r}}(\{1\}).$ By 8.1 , $G$ is an induced subgraph of $G_{F_{d}}(\{1\})$ as well.
+Since $q^{m_{i}}>n$, we know that $G_{n}$ is an induced subgraph of $G_{F_{r}}(\{1\}).$ By 8.1 , $G$ is an induced subgraph of $G_{F_{d}}(\{1\})$ as well. ∎
+We obtain the following result which says in particular that every graph is an induced subgraph of a gcd-graph.
 Corollary 8.3.
@@
 $\{f_{1},f_{2},\ldots,f_{k}\}=\{h_{1},h_{2},\ldots,h_{k}\}.$
-Let $I$ be the ideal generated by $h_{0}$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$ By 5.4 the induced graph on $I$ is naturally isomorphic to the unitary Cayley graph $G_{h}(\{1\})$.
+Let $I$ be the ideal generated by $h_{0}$ in $\operatorname{\mathbb{F}}_{q}[x]/(f).$ By 5.4 the induced graph on $I$ is naturally isomorphic to the unitary Cayley graph $G_{h}(\{1\})$.
 This shows that $G$ is an induced subgraph of $G_{f}(D).$
@@
   2. (2)
-    $G$ is an induced subgraph of $G_{f}(D)$ where $D=\{f_{1},f_{2},\ldots,f_{k}\}$.
+    $G$ is an induced subgraph of $G_{f}(D)$ where $D=\{f_{1},f_{2},\ldots,f_{k}\}$?
 In general, the answer is no.
 There seem to be some subtle constraints.
-We discuss here a particle one.
+We discuss here a particular one.
 Let $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})$ and assume further that $g\neq f_{i}$ for all $1\leq i\leq k$.
 If such $f$ exists, then $g\mid f$ and there is a canonical map
-$\Phi:\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}[x]/g.$
-If we take a subset $S\subset\operatorname{\mathbb{F}}_{q}[x]/f$ such that $|S|>|\operatorname{\mathbb{F}}_{q}[x]/g|$, then there exists $a,b\in S$ such that $a\neq b$ and $\Phi(a)=\Phi(b).$ Consequently, $\Phi(a-b)=0$ or $g\mid a-b.$ By definition, $(a,b)\not\in E(G_{f}(D)).$ Consequently, we have the following upper bound for the clique number of $G_{f}(D)$
-$\omega(G_{f}(D))\leq|\operatorname{\mathbb{F}}_{q}[x]/g|.$
-This shows that if $\omega(G)>|\operatorname{\mathbb{F}}_{q}[x]/g|$ then $G$ cannot be an induced subgraph of $G_{f}(D).$
+$\Phi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}[x]/(g).$
+If we take a subset $S\subset\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $|S|>|\operatorname{\mathbb{F}}_{q}[x]/(g)|$, then there exists $a,b\in S$ such that $a\neq b$ and $\Phi(a)=\Phi(b).$ Consequently, $\Phi(a-b)=0$ or $g\mid a-b.$ By definition, $(a,b)\not\in E(G_{f}(D)).$ Consequently, we have the following upper bound for the clique number of $G_{f}(D)$
+$\omega(G_{f}(D))\leq|\operatorname{\mathbb{F}}_{q}[x]/(g)|.$
+This shows that if $\omega(G)>|\operatorname{\mathbb{F}}_{q}[x]/(g)|$ then $G$ cannot be an induced subgraph of $G_{f}(D).$
 We can overcome the constraint by enlarging the base field $\operatorname{\mathbb{F}}_{q}.$
@@
 Let $m$ be a positive integer such that $q^{m}>n$ and $g=\textrm{lcm}(f_{1},f_{2},\ldots,f_{k})\in\operatorname{\mathbb{F}}_{q}[x].$ By Galois theory, there exists a polynomial $h\in\operatorname{\mathbb{F}}_{q}[x]$ with at least $r$ distinct irreducible factors and $\gcd(h,g)=1.$ Let $f=hg$.
-Let $I$ be the ideal in $\operatorname{\mathbb{F}}_{q^{m}}[x]/f$ generated by $f_{1}.$ Then, by 5.4, the induced subgraph on $I$ is isomorphic to the unitary Cayley graph $G_{\bar{f}}(\{1\})$ where $\bar{f}=f/f_{1}.$ Suppose that $\bar{f}=g_{1}^{a_{1}}g_{2}^{a_{2}}\ldots g_{t}^{a_{t}}$ be the factorization of $f/f_{1}$ over $\operatorname{\mathbb{F}}_{q^{m}}[x].$ Then, by the choice of $f$, $t\geq r.$ By the Chinese remainder theorem, we know that $\operatorname{\mathbb{F}}_{q^{m}}^{t}$ is a subring of $\operatorname{\mathbb{F}}_{q^{m}}[x]/\bar{f}.$ Therefore, $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}=\prod_{i=1}^{t}K_{q^{m}}$ is an induced subgraph of $G_{\bar{f}}(\{1\})$.
+Let $I$ be the ideal in $\operatorname{\mathbb{F}}_{q^{m}}[x]/(f)$ generated by $f_{1}.$ Then, by 5.4, the induced subgraph on $I$ is isomorphic to the unitary Cayley graph $G_{\bar{f}}(\{1\})$ where $\bar{f}=f/f_{1}.$ Suppose that $\bar{f}=g_{1}^{a_{1}}g_{2}^{a_{2}}\ldots g_{t}^{a_{t}}$ be the factorization of $f/f_{1}$ over $\operatorname{\mathbb{F}}_{q^{m}}[x].$ Then, by the choice of $f$, $t\geq r.$ By the Chinese remainder theorem, we know that $\operatorname{\mathbb{F}}_{q^{m}}^{t}$ is a subring of $\operatorname{\mathbb{F}}_{q^{m}}[x]/\bar{f}.$ Therefore, $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}=\prod_{i=1}^{t}K_{q^{m}}$ is an induced subgraph of $G_{\bar{f}}(\{1\})$.
 Since $t\geq r$ and $q^{m}\geq n$, by 8.1 $G_{n}$ is an induced subgraph of $G_{\operatorname{\mathbb{F}}_{q^{m}}^{t}}$.
@@
 We thank Professor Torsten Sander for the enlightening correspondence about gcd-graphs and related arithmetic graphs.
-We are grateful to Professor Emanuel Kolwaski for some helpful comments on Fekete polynomials and for writing the book [19].
+We are grateful to Professor Emmanuel Kolwaski for some helpful comments on Fekete polynomials and for writing the book [20].
 We thank Professor Ki-Bong Nam for his interest in this topic and his feedback on the first draft.
 Finally, we thank Minh-Tam Trinh for his discussions, encouragement, and support at the initial stage of this work.
+We are also grateful to the referee for their comments and valuable suggestions which we have used to improve our exposition.